### v3

# with Mixture of Frequency Adaptation

Abstract
Parameter-Efficient Fine-Tuning (PEFT) is a technique that allows us to adapt powerful Foundation Models (FMs) to diverse downstream tasks while preserving and unleashing their inherent capabilities. However, we have observed that existing PEFT methods, which are often designed with natural imagery in mind, struggle when applied to Remote Sensing (RS) scenarios. This is primarily due to their inability to handle artifact influences, a problem particularly severe in RS image features. To tackle this challenge, we introduce Earth-Adapter, the first PEFT method specifically designed for RS artifacts conquering. Earth-Adapter introduces a novel Mixture of Frequency Adaptation process that combines a Mixture of Adapter (MoA) with Discrete Fourier Transformation (DFT). By utilizing DFT, Earth-Adapter can decompose features into different frequency components, precisely separating artifacts from original features. The MoA then dynamically assigns weights to each adapter expert, allowing for the combination of features across various frequency domains. These simple-yet-effective approaches enable Earth-Adapter to more efficiently overcome the disturbances caused by artifacts than previous PEFT methods, significantly enhancing the FMs’ performance on RS scenarios. Experiments on Domain Adaptation (DA), and Domain Generalization (DG) semantic segmentation benchmarks showcase the Earth-Adapter’s effectiveness. Compared with baseline Rein, Earth-Adapter significantly improves 9.0% mIoU in DA and 3.1% mIoU in DG benchmarks. Our code will be released at Earth-Adapter.
1 Introduction
Tremendous last year breakthroughs in Foundation Models (FMs) [14, 45, 54, 40, 25, 55, 22, 3] have ignited transformative progress in Computer Vision (CV), demonstrating unprecedented capabilities across both high-level and low-level vision tasks. Based on these FMs, considerable Remote Sensing (RS) research groups have also constructed large-scale, state-of-the-art (SOTA), and impressive Geospatial Foundation Models (GFMs) [5, 51, 21, 65, 70, 43, 36, 23, 50, 38], significantly promoting the development of RS imagery interpretation field. However, when adapting these FMs to downstream tasks, how to efficiently keep and unleash the power of FMs becomes a key problem. Under this context, Parameter-Efficient Fine-Tuning (PEFT) methods [31, 37, 28, 8, 20] become the pivotal method due to their superior parameter-performance trade-off.
Existing PEFT methods designed for language or nature imagery have made great progress with excellent works, such as LoRA [31], Visual Prompt Tuning (VPT) [37]. However, most of them show performance drops in RS cross-domain experiments. We argue the main reason lies in the influences of artifacts in the features. As shown in Figure 1 (a), we show the PCA visualization of features of DINOv2-L, which contains obvious redundant artifacts. Although a part of existing PEFT methods also are shown to be capable of handling artifacts [12], they still underestimate the artifact disturbances in RS imagery. The difference in the distribution of artifacts between natural and RS images is significant. In natural images, artifacts typically surround foreground objects [12], such as humans or animals, and the disturbances they cause are relatively limited. In contrast, RS images, due to their overhead perspective, lack centralized subjects and contain multiple coexisting multi-scale targets. For example, a single RS image may simultaneously include large-scale agricultural regions and fragmented road networks. As a result, artifacts are almost situated everywhere in RS images shown in Figure 1 (a), causing severe interference to domain-invariant representation learning, which is pivotal for cross-domain scenarios.
[…]
In our view, to address the influence of artifacts, one of the core ideas is to ‘divide-and-conquer’. Drawing on the principle that high-frequency (HF) signals encode local details (e.g., edge, contour) while low-frequency (LF) signals encode global structure (e.g., object surface), we leverage Discrete Fourier Transformation (DFT) as the ‘divide’ step. As illustrated in Figure 1 (b), we successfully leverage the DFT effectively to isolate artifacts into HF signals and splits coherent global features into LF signals. Subsequently, to ‘Conquer’ artifact-related issues, we adopt a Mixture of Adapter (MoA) architecture for PEFT design. The MoA router dynamically assigns weights to adapter experts, selecting the most relevant features for processing in each iteration. In this way, all features are aggregated into optimized representations, mitigating artifacts disturbances.
Based on our motivation and findings, we further present Earth-Adapter, a simple-yet-effective RS-specialized PEFT method designed for RS artifact conquering. Specifically, Earth-Adapter consists of a novel Mixture of Frequency Adapter architecture consisting of Mixture of Adapter (MoA) with Discrete Fourier Transformation (DFT). Earth-Adapter first disentangles geospatial features by DFT, and leverages multiple experts to conquer artifacts by adaptively aggregating and learning features with suitable frequency. In this fashion, Earth-Adapter can effectively alleviate the artifacts’ disturbances for robust domain-invariant feature learning. Then gradually introduces FMs to cross the geospatial gaps from the frequency perspective.
In experiments, we meticulously test Earth-Adapter on Domain Adaptation (DA), and Domain Generalization (DG) semantic segmentation tasks across five widely-used datasets. Comparing with existing PEFT methods, Earth-Adapter achieves State-Of-The-Art (SOTA) performance on DA and DG benchmarks. In DA and DG benchmarks, Earth-Adapter outperforms our baseline Rein with 8.9% and 3.2% mIoU, respectively. Additionally, we also conduct in-depth analysis of the potential of Earth-Adapter. The core contributions can be summarized as follows:
  • We propose a novel perspective to alleviate artifact-disturbance problems in RS imagery by frequency-guided divide-and-conquer strategy.
  • We introduce a simple-yet-effective PEFT method, Earth-Adapter. It is also the first RS-specialized PEFT approach designed for cross-domain semantic segmentation tasks.
  • Extensive comparisons validate the SOTA performance of Earth-Adapter on both DA and DG benchmarks. More ablation studies also demonstrate the high efficiency, robustness, and explainability of Earth-Adapter.
2 Related Work
2.1 RS Cross-Domain Semantic Segmentation
As a classical and fundamental task in the realm of computer vision, semantic segmentation [46, 57] not merely holds profound practical significance, but also provides an interface for researchers to dissect the high-level semantic representation of images. With the growing need for applying models to out-of-domain data, cross-domain segmentation, including domain adaptation (DA) [29, 30, 19, 41, 33] and domain generalization (DG) [67, 21], has gained prominence. DA and DG are the two main paradigms, where DA aims to adapt a model from a labeled source domain to an unlabeled target domain, while DG focuses on training a model on only the source domain to generalize to unseen target domains. In the RS field, cross-domain task is crucial due to the diversity of data, and lots of DA methods have been applied to improve the generalization ability of segmentation models, with applications in urban planning [56], land resource management [60], and environmental protection [58]. Compared with DA, the number of DG research [34, 42] in RS is much less than DA. Although these works [72, 35, 16] make contributions, most of them only focus on specialized models rather than leveraging the power of FMs, leading to limited cross-domain capabilities.
2.2 Geospatial Foundation Models
The emergence of Large Language Models (LLMs) [4] has prompted a shift in computer vision, leading to the development of Multimodal Large Language Models (MLLMs) [48, 54, 44, 9] and Vision Foundation Models (VFMs) [45, 52, 24, 40]. FMs [3] in RS called GFMs [64, 26, 63, 62, 6] also can be divided into MLLMs and VFMs. Among MLLMs, RemoteCLIP [43] is the first exploration, supporting classification, retrieval, and object counting through a contrastive learning framework based on CLIP [54]. MetaEarth [70] advances diffusion models to generate multi-resolution RS images for any specified region based on both image and text conditions. Skysense [23] captures spatiotemporal information and performs well across diverse scenarios. Although these FMs obtain strong power by large-scale pertaining, how to effectively unleash their power to diverse downstream tasks is still a challenge.
2.3 PEFT in RS
PEFT approaches, such as adapter [28, 27, 31] are proposed to fine-tune large pre-trained models to downstream tasks in Natural Language Processing (NLP) by introducing light-weight learnable parameters. Subsequently, the PEFT has begun to be transferred to vision[69, 1, 39] tasks and rapidly become the focus of researchers. Visual Prompt Tuning (VPT) [37] first introduced ‘prompt tuning’ into the vision area as learnable vectors. The thought of this paradigm is also similar to the adversarial re-programming [15]. Currently, cross-domain works also focus on PEFT, such as [18] which is the first prompt tuning method in Domain Adaptation (DA) [59, 17, 20] and [67] is the first work leveraging PEFT to fine-tuning VFMs for Domain Generalization (DG) [2]. In the RS field, the exploration of PEFT is also vigorous. For example, UPetu [13] addresses storage for dense prediction tasks. TEA [32] uses an adapter network and top-down guidance for detection and segmentation tasks. [53] also leverage LoRA to solve oriented detection task. However, for the cross-domain semantic segmentation tasks, the research of PEFT is relatively blank, and existing methods can not handle the influence of artifacts. Earth-Adapter is the first PEFT for this area and can effectively tackle the artifacts.
3 Method
3.1 Preliminary
We first denote the input RS images as $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, the corresponding label of $\mathbf{x}$ as $\mathbf{y}\in\mathbb{R}^{H\times W\times K}$, and the semantic segmentation prediction of $\mathbf{x}$ as $\mathbf{y}^{\prime}\in\mathbb{R}^{H\times W\times K}$, where $K$ represents the number of semantic classes. Here, we’ll illustrate our innovative Earth-Adapter overview.
The proposed Earth-Adapter is composed of two key components: the DFT and MoA. MoA consists of a Spatial Adapter, a LF Adapter, a HF Adapter, and a router. We employ $R_{\xi}$, parameterized by $\xi$, to denote the router and $E^{i}_{\epsilon},\;i\in\{1,2,3\}$, parameterized by $\epsilon$, to represent the three adapter experts. For the network architecture, we adopt the framework introduced by Rein [67], integrating DINOv2 [52] as the backbone, represented as $f_{\phi}$ and parameterized by $\phi$, alongside Mask2Former [10] as the decoder, denoted by $f_{\theta}$ and parameterized by $\theta$. Section 3.2 illustrates the Earth-Adapter’s design, while Section 3.3 elaborates on our cross-domain training framework.
3.2 Details of Earth-Adapter
The core idea of Earth-Adapter is the seamless integration of highly efficient, learnable layers with minimal parameters into the frozen FMs.
Optimization objective. Before employing the Earth-Adapter, the optimization objective of fine-tuning FMs is to identify a set of parameters that minimize the loss of the entire model on the downstream task:
$\mathop{\arg\min}_{\theta}\,\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\mathcal{L}_{\text{seg}}\left(f_{\theta}(f_{\phi}(\mathbf{x})),\mathbf{y}\right),$ (1)
where $\mathcal{D}$ is the dataset, $f_{\phi}\cdot f_{\theta}$ is the segmentation model, and $\mathcal{L}_{seg}$ is the loss function, here we use the default CE loss. After activating our Earth-Adapter, the optimization objective becomes:
$\mathop{\arg\min}_{\theta,\epsilon,\xi}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\mathcal{L}_{\text{seg}}\left(f_{\theta}\left(\underbrace{R_{\xi}\circ E_{\epsilon}}_{\text{Earth-Adapter}}(f_{\phi}^{*}(\mathbf{x}))\right),\mathbf{y}\right).$ (2)
During training, the original parameters of the backbone are kept frozen (denoted as $f_{\phi}^{*}$), ensuring they remain unaltered. In the following description, we denote the segmentation network combined with Earth-Adapter as $G_{\theta,\xi,\epsilon}$.
Parallel Frequency Adaptation. Let $\mathbf{F}_{i}\in\mathbb{R}^{n\times c}$ denote the visual feature at the $i$-th layer of the backbone, where $n$ represents the token sequence length and $c$ denotes the channel dimension. Our frequency adaptation operates through parallel processing streams: The first Spatial Adapter employs a low-rank projection to refine the spatial feature:
$\Delta\mathbf{F}^{spatial}_{i}=\text{Adapter}_{1}^{i}(\mathbf{F}_{i}^{T}),$ (3)
where $\Delta\mathbf{F}^{spatial}_{i}$ means the processed i-th layer Spatial-Domain features, and $\textbf{Adapter}_{1}$ represents a nonlinear mapping layer composed of two low-rank matrices:
$\textbf{Adapter}_{i}=\textbf{W}_{up}(\text{Relu}(\textbf{W}_{down}(\cdot))).$ (4)
The Frequency Adapter decompose features through spectral analysis. We first reshape $\mathbf{F}^{spatial}$ to $({C,H,W})$ and apply DFT on spatial features as $\mathcal{FT}(\mathbf{F}_{spatial})$. When splitting the frequency domains, we employ a fixed frequency cutoff $\rho$ to decompose the spectrum into high and low frequency components. Subsequently, these components are transformed back into features via the Inverse Fourier Transform (IFT), yielding LF and HF features:
$\mathbf{F}^{low}_{i}=\mathcal{FT}^{-1}(\mathbf{M}\odot\mathcal{FT}(\mathbf{F}_{spatial})),$ (5)
$\mathbf{F}^{high}_{i}=\mathcal{FT}^{-1}((1-\mathbf{M})\odot\mathcal{FT}(\mathbf{F}_{spatial})).$ (6)
The frequency mask $\mathbf{M}\in\{0,1\}^{H\times W}$ is defined by:
$\mathbf{M}(u,v)=\begin{cases}1&,\thinspace\text{if }\max(|u-\frac{H}{2}|,|v-\frac{W}{2}|)\leq\rho\frac{H}{2}\\ 0&,\thinspace\text{otherwise}\end{cases}$ (7)
Thereafter, the LF and HFfeatures are independently passed through two distinct low-rank linear projection layers, generating frequency adaptation adjustments:
$\Delta\mathbf{F}^{low}_{i}=\text{Adapter}_{2}^{i}(\mathbf{F}^{low}_{i}),$ (8)
$\Delta\mathbf{F}^{high}_{i}=\text{Adapter}_{3}^{i}(\mathbf{F}^{high}_{i}).$ (9)
Mixture of Adapters. We implement dynamic feature fusion through a Router that learns optimal combinations of feature adjustment according to the original visual features. The Router weight is computed by channel-wise attention:
$\mathbf{w}_{i}=\text{Softmax}(R_{\xi}(\mathbf{F}_{i})),$ (10)
where $\mathbf{w}_{i}$ represents weights for spatial, LF, and HF components. The final feature adjustment is calculated as:
$\Delta\mathbf{F}_{i}=\alpha_{i}\sum_{k=1}^{3}\mathbf{w}^{(k)}\Delta\mathbf{F}^{(k)}_{i},$ (11)
where $\alpha_{i}$ is a learnable scaling parameter with small initial value, and $k\in\{spatial,low,high\}$. The complete feature optimization also maintains residual structure:
$\bar{\mathbf{F}_{i}}=\mathbf{F}_{i}+\Delta\mathbf{F}_{i}.$ (12)
3.3 Training FrameWork
In this study, we explore the application of Earth-Adapter across two distinct tasks: DG and UDA Semantic Segmentation. For DG Semantic Segmentation tasks, we train Earth-Adapter with Mask2former in end-to-end supervised learning as Equation 2. For UDA Semantic Segmentation, we use the self-training framework introduced in DACS [61] without using target domain labels. The pseudo labels for target domain images are produced by a teacher network $G_{\theta,\xi,\epsilon}^{\dagger}$:
$\mathbf{\bar{y}_{T}}^{(j,k)}=[k=\mathop{\arg\max}_{k^{\prime}}G_{\theta,\xi,\epsilon}^{\dagger}(\mathbf{x_{T}})))^{(j,k^{\prime})}],$ (13)
where []delimited-[][$\cdot$][ ⋅ ] denotes the Iverson bracket, and $j\in HW$ is the spatial index of the pixel in the image. The teacher network remains inactive during the training process, with its parameters being updated based on an exponential moving average (EMA) of the student model’s parameters:
$G_{\theta,\xi,\epsilon,t+1}^{\dagger}=\alpha G_{\theta,\xi,\epsilon,t}^{\dagger}+(1-\alpha)G_{\theta,\xi,\epsilon,t}$ (14)
Here $t$ denotes the $t_{th}$ training iteration. The hyperparameter $\alpha$ controls the update speed of the teacher network’s parameters. After obtaining the pseudo-labels, we perform category-wise mixing of samples from the source and target domains:
$\mathbf{x}_{mix}$ $=\mathbf{x_{S}}\odot\mathcal{M}+\mathbf{x_{T}}\odot(1-\mathcal{M})$ (15)
$\mathbf{y}_{mix}$ $=\mathbf{y_{S}}\odot\mathcal{M}+\mathbf{\bar{y}_{T}}\odot(1-\mathcal{M})$
where $(\mathbf{x_{S}},\mathbf{y_{S}})$ is the data sample from the source domain, and $\mathcal{M}$ is generated by randomly selecting half of the categories based on the labels of the source domain $\mathbf{y_{S}}$. The Optimization objective for the UDA task if defined by:
$\mathop{\arg\min}_{\theta,\epsilon,\xi}$ $\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}\cup\mathcal{D}_{mix}}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}),\mathbf{y}\right)$ (16)
$\quad+\lambda_{uda}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}_{mix}),\mathbf{y}_{mix}\right).$
$\lambda_{uda}$ is the weight parameter for UDA loss.
4 Experiments
Backbone Methods Cross-Domain Adaptation Avg. Cross-Domain Generalization Avg.
P2V V2P R2U U2R P2V V2P R2U U2R
MiT-B5 [68] DAFormer [29] 64.4 54.8 52.7 42.5 53.6 42.4 41.4 54.2 39.9 44.5
HRDA [30] 67.6 58.6 53.2 35.3 53.7 33.1 31.1 54.2 39.8 39.6
DINOv2-L [52] Frozen 21.0 7.2 21.5 11.8 15.4 57.9 49.4 57.1 42.7 51.8
Full Fine-Tune 11.3 16.5 23.1 10.6 15.4 12.6 19.6 33.1 19.7 21.3
Adapter [28] 66.4 59.3 55.9 46.2 57.0 56.0 47.4 57.6 41.8 50.7
LoRA [31] 17.8 18.8 24.5 26.0 15.7 20.3 25.6 29.3 21.9 24.3
VPT [37] 66.2 59.3 55.3 48.0 57.2 59.2 52.3 57.4 44.9 53.5
AdaptFormer [7] 12.9 15.3 25.1 22.2 18.9 14.0 19.3 31.0 15.8 20.0
Rein (Baseline) [67] 60.2 60.9 52.8 26.0 50.0 60.8 52.5 55.8 43.4 53.1
Earth-Adapter (Ours) 67.7 (+7.5) 62.2 (+1.3) 55.9 (+3.1) 50.0 (+24.0) 59.0 (+9.0) 64.9 (+4.1) 55.1 (+2.6) 59.0 (+3.2) 45.7 (+2.3) 56.2 (+3.1)
Improvements over the second best +0.1 +1.3 +0.0 +2.0 +1.8 +4.1 +2.5 +1.4 +0.8 +2.7
Table 1: Performance Comparison between representative DA models, PEFT methods, and our Earth-Adapter on DA and DG benchmarks. The Bold font in the table means the SOTA performance. The ‘+’ in the bracket means the improvement compared with our baseline Rein. The last line of data means the improvement compared with the second-best performance of all other methods.
Methods Setting Surf Bldg Vegt Tree Car Clut mIoU (%) Setting Bkgd Bldg Road Water Barr Frst Agri mIoU (%)
Frozen P2V (DA) 38.6 24.1 11.9 47.2 2.5 1.5 21.0 U2R (DA) 42.6 17.2 6.7 9.6 0.7 5.7 0.0 11.8
Rein [67] 75.4 88.6 50.3 73.4 59.0 14.5 60.2 27.9 41.9 29.3 51.2 0.8 53.2 0.0 29.2
Earth-Adapter (Ours) 80.3 89.5 66.3 74.4 61.9 33.6 67.7 56.3 63.4 48.7 66.6 13.7 38.4 63.2 50.0
Frozen P2V (DG) 70.6 79.9 46.1 72.1 58.6 20.4 57.9 U2R (DG) 56.7 54.8 41.3 67.0 14.6 15.1 49.2 42.7
Rein [67] 74.4 85.1 51.0 23.0 57.1 24.5 60.8 53.0 57.8 41.2 61.1 9.4 23.4 57.8 43.4
Earth-Adapter (Ours) 76.3 85.4 59.0 74.1 54.2 40.6 64.9 52.6 55.7 44.1 64.6 17.0 29.6 56.1 46.7
Table 2: Performance Comparison of Per-Class between Frozen DINOv2-L, Rein, and Earth-Adapter on Potsdam RGB-to-Vaihingen. Both Rein and ours are based on the Frozen DINOv2-L. The Bold font means the best performance.
Backbone Methods Params P2V(DG) V2P(DG) Avg.
CLIP [54] (ViT-Large) Frozen 0.0M 43.7 44.4 44.1
Full Fine-Tune 304.2M 11.4 20.3 15.9
Earth-Adapter 2.4M 49.7 50.2 50.0 (+5.9)
DINOv2 [52] (Small) Frozen 0.0M 44.3 43.3 43.8
Full Fine-Tune 22.1M 41.9 41.8 41.9
Earth-Adapter 2.4M 47.5 44.8 46.2 (+2.4)
DINOv2 [52] (Base) Frozen 0.0M 51.4 49.5 50.5
Full Fine-Tune 86.6M 46.7 49.8 49.8
Earth-Adapter 2.4M 53.0 51.8 52.4 (+1.9)
DINOv2 [52] (Large) Frozen 0.0M 57.9 49.4 53.7
Full Fine-Tune 304.2M 12.6 19.6 16.1
Earth-Adapter 2.4M 64.9 55.1 60.0 (+6.3)
Table 3: Ablation studies of Earth-Adapter on different foundation backbones. Both P2V and V2P are the DG benchmarks. Results validate that Earth-Adapter’s effectiveness on diverse backbone networks.
Adapter + HF + LF P2V (DG) V2P (DG) Avg.
1 - - 61.0 52.8 56.9 (+0.0)
- 1 - 59.4 50.9 50.9 (-6.0)
- - 1 59.1 51.5 55.3 (-1.6)
2 - - 61.6 51.4 56.5 (-0.4)
3 - - 64.0 52.7 58.4 (+1.5)
4 - - 61.4 53.3 57.4 (+0.5)
5 - - 62.5 52.5 57.5 (+0.6)
1 1 1 64.9 55.1 60.0 (+3.1)
2 1 1 61.3 54.7 58.0 (+1.1)
3 1 1 61.6 54.6 58.1 (+1.2)
Table 4: Ablation of Adapters’ Combination. The ‘Adapter’ means the number of vanilla Adapter that accepts Spatial features. The ‘+ HF’ and ‘+ LF’ represent the number of adapters that accept high and low-frequency features as input. The ‘+’ in the bracket represents the gain compared with one Adapter.
Backbone Methods Params $\downarrow$ Train Speed (s/iter) $\downarrow$ Infer Speed (s/iter) $\downarrow$ mIoU (%) $\uparrow$
DINOv2-L Full 304.2M 0.80 0.11 15.4
Rein 3.0M 0.68 0.12 50.0
Earth-Adapter 2.3M(1.0Msimilar-$\sim$3.9M) 0.71 0.12 59.0 (+9.0)
Table 5: Comparison on Parameters, Speed, and Performance between FFT, Rein, and Earth-Adapter. All results are the average score of DA benchmarks, demonstrating the excellent trade-off of Earth-Adapter.
Figure 2: Visualization of Predicted Segmentation Maps We Compare Earth-Adapter with the Frozen DINOv2-L backbone and our baseline Rein on eight cross-domain benchmarks. For the Potsdam and Vaihingen color map, white is the Impervious surface, red is the clutter, blue is the building, color is the low vegetation, green is the tree, and yellow is the car. For LoveDA color map, red is the building, yellow is the road, blue is the water, purple is the barren, green is the forest, brown is the agriculture.
4.1 Experimental setup
Datasets and Benchmark. We conduct all experiments on several widely used remote sensing image segmentation datasets: Potsdam [49] and Vaihingen [49], both created by the International Society for Photogrammetry and Remote Sensing (ISPRS) and LoveDA [66]. The Potsdam dataset includes 38 very high resolution (VHR) True Orthophotos (TOP) images with 6,000 × 6,000 pixels. The Potsdam dataset has different imaging modes including IR-R-G channels, R-G-B channels, and R-G-B-IR channels. We use the RGB band for the following experiments. The Vaihingen dataset consists of 33 VHR TOP images with an average size of 2,494 ×2,064 pixels. The LoveDA dataset comprises a total of 4,190 images, each with a resolution of 1,024×1,024 pixels. Based on the aforementioned dataset, we follow MMSegmentation settings to split the train and val sets, crop images, and establish four benchmarks including both DA and DG: Potsdam to Vaihingen (P2V), Vaihingen to Potsdam (V2P), Rural to Urban (R2U), Urban to Rural (U2R). In our experimental setup, all images are cropped to a size of $512\times 512$.
Implementation Details. We use MMSegmentation [11] as our training and evaluation framework. We employ Mask2former [10] as our decoder, which is a highly efficient and effective algorithm for semantic segmentation tasks. For backbone, we utilize the Dinov2 [52] series models to extract features, which are subsequently utilized as input for the Mask2former. During training, we utilize AdamW [47] as the optimizer, with a learning rate of 1e-5 for the backbone, 1e-4 for the decoder, and 1e-4 for the relevant parameters in PEFT. Within the DACS training framework, we set $\alpha$ to 0.99 and $\lambda_{uda}$ to 0.5.
4.2 Cross-Domain Performance Comparison
As depicted in Table 1, we conduct a comparative analysis of the existing mainstream PEFT methods, such as Adapter [28], LoRA [31], and VPT [37], against our Earth-Adapter across eight cross-domain benchmarks,with state-of-the-art method Rein [67] in natural scene as our baseline. These benchmarks encompass four DA scenarios (P2V, V2P, R2U, and U2R) and four DG scenarios that mirror the DA ones.
Figure 3: Visualization and PCA of Adapters’ Feature Maps. ‘Agg. Feature’ represents the aggregated adapters’ features. ‘PCA’ represents the Principal Component Analysis of features. All visualizations represent feature maps, not heatmaps. Thus only the semantic boundaries within the features should be focused rather than color intensities.
Figure 4: Parameter-sensitive experiments of the cut-off frequency of DFT. The results on P2V (DA) and U2R (DG) benchmarks show Earth-Adapter’s robustness to the cut-off frequency.
Starting with the DA experiments, when compared to the performance of DA-specialized models, a significant number of existing PEFT methods exhibit performance collapse. For instance, LoRA achieves only 17.8% mIoU in the P2V (DA) benchmark when handling RS (Remote Sensing) images, a figure even lower than that of the Frozen backbone DINOv2-L. A similar trend is observed in Full Fine-Tuning, AdaptFormer, and Rein. We posit that this issue may stem from the DA self-training paradigm, which involves intensive image-mixing operations that these PEFT methods struggle to adapt to. However, two PEFT methods, Adapter and Rein, demonstrate relatively robust performance in the DA setting. When compared to these methods, our Earth-Adapter proves to be more effective in adapting to target domains. It outperforms baseline Rein by 7.5% mIoU in P2V, 3.1% mIoU in R2U, 24.0% mIoU in U2R, and 9.0% mIoU on average across the DA benchmarks. Notably, in the U2R (DA) benchmark, Earth-Adapter significantly mitigates the collapse of Rein, boosting the mIoU from 26.0% to 50.0%.
Shifting to the DG experiments, the performance collapse persists, though the trend is somewhat alleviated. In contrast to DA, the Frozen model reveals the potential generalizability of DINOv2-L. Nevertheless, FFT, LoRA, and AdaptFormer still exhibit limitations. On the other hand VPT, and Rein continue to show strong transfer capabilities, achieving 53.5% and 53.1% mIoU on average, surpassing the Frozen model by 1.7%, and 1.3% mIoU. Similar to the DA benchmark, Earth-Adapter also achieves state-of-the-art (SOTA) performance in DG, with improvements of 3.1% and 4.4% mIoU over the baseline Rein and Adapter, respectively.
4.3 Ablation Study
Analysis on Per-class Performance. In this part, we demonstrate the detailed performance of each class on four benchmarks: P2V (DA), P2V (DG), U2R (DA), and U2R (DG) between Frozen DINOv2-L, Rein, and our Earth-Adapter. All the results are shown in Table 2, and we first discuss the P2V experiment results.
In the P2V (DA), both Rein and Earth-Adapter show significant improvement on the Vegetation class compared with the Frozen DINOv2-L. Nevertheless, Earth-Adapter achieves a much higher performance than Rein, with 7.5% mIoU. A similar trend can be observed in the P2V (DG), where Earth-Adapter surpasses the other two methods by 7.0% and 4.1% mIoU respectively. Looking at the performance changes of the Vegetation and Tree classes, something interesting stands out. In the P2V (DA), both Earth-Adapter and Rein perform better on the Tree class than the Frozen. In contrast, in the P2V (DG), the performance of the Frozen with Rein on the Tree class drops drastically from 74.1% to 23.0% mIoU. As indicated in prior research [21], Rein’s features struggle to distinguish between the Vegetation and Tree classes due to their similar appearance from the RS perspective, especially in DG setting which faces more unpredictable domain shifts than DA. Moreover, in P2V (DG), comparing Rein with the Frozen, almost all other classes improve except for the Tree class. Thus, it can be empirically inferred that the improvement of the Vegetation class by Rein comes at the cost of a decline in the performance of the Tree class. However, Earth-Adapter can reduce this negative impact. It does this by optimizing the representation to tell the two classes apart and boost their performance.
Figure 5: Parameter-sensitive experiments of the projector dimension in Adapter. The results on P2V (DA) and U2R (DG) benchmarks show Earth-Adapter’s robustness to dimensions.
In the LoveDA experiments, as previously discussed, both the Frozen and Rein methods suffer from performance collapse in the U2R (DA) setting. When examining the performance at a more class level, it becomes evident that the primary decline for Rein occurs in the Barren and Agriculture classes. The performance drop in the Barren class can be attributed to it being the most sparse class within the LoveDA dataset. However, the performance of both the Frozen and Rein methods on the Agriculture class is a concerning 0.0% mIoU. In our opinion, the main factor probably is the presence of artifacts, as illustrated in Figure 1. Most PEFT methods from the CV field are unable to effectively handle the disturbances caused by these artifacts. Nevertheless, Earth-Adapter is capable of effectively utilizing the target domain data, achieving a notable 13.7% mIoU in the Barren class. Furthermore, as depicted in Figure 1, Earth-Adapter has a significant advantage with its frequency-domain divide-and-conquering ability, which enables it to improve the Agriculture performance from 0 to 63.2% mIoU. In the U2R (DG) experiments, both the Frozen and Rein methods demonstrate their effectiveness and significantly address the issues present in the DA setting. Additionally, Earth-Adapter achieves SOTA performance with improvements across most classes.
Performance on Different Backbones. Besides the DINOv2-L backbone in Table 1, Earth-Adapter is also applied to CLIP (ViT-L) and other DINOv2 scales: ViT-Small and ViT-Base, with results in Table 3. These results show Earth-Adapter improves performance across diverse backbones with varying parameter scales. Notably, Earth-Adapter achieves larger improvements on larger models, such as ViT-L, with 7.0% and 5.7% mIoU gains, more than those on small and base scale ViTs. This suggests Earth-Adapter has a greater performance upper bound.
Prediction Visualization Analysis. In Figure 2, we visualize the predicted segmentation map of the Frozen DINOv2-L, Rein, and Earth-Adapter. It is worth noting that in the U2R (DG) example, Rein’s prediction is worse than the original backbone, which means Rein can not adapt to RS image well, which leads to a negative effect on the backbone features. In contrast, on the backbone’s basement, the Earth-Adapter keeps the good details of backbone features, and further optimizes the representation of agriculture class. And in other experiments, Earth-Adapter also exhibits better prediction than Rein and Frozen backbone, showing a higher performance upperbound.
Benchmark Adapter dim Cutoff frequency Activation layer mIoU(%)
P2V (DA) 64 0.3 [0,1,2] 67.7
P2V (DG) 24 0.3 [18,19,20,21,22,23] 64.9
V2P (DA) 32 0.2 [21,22,23] 62.2
V2P (DG) 64 0.3 [18,19,20,21,22,23] 55.1
R2U (DA) 16 0.3 [21,22,23] 55.9
R2U (DG) 64 0.3 [18,19,20,21,22,23] 59.0
U2R (DA) 32 0.2 [18,19,20,21,22,23] 50.0
U2R (DG) 64 0.3 [21,22,23] 45.7
Table 6: Hyperparameter Configuration of dimension, cutoff frequency, and activation layer of Frequency Adapters in Earth-Adapter. Notably, Spatial Adapter is activated in all backbone layers. All the results are the best performance.
Feature Visualization Analysis. In Figure 3, we visualize features captured by three adapters on P2V (DG) and R2U (DG) benchmarks. As discussed in the introduction, LF features focus on coarse-grained and global semantics, while HF features show detailed representations. The aggregated features are a weighted summation of the three different frequency features. PCA visualizations clearly show each feature’s characteristics. Take P2V in the figure as an example, artifacts are almost filtered into high-frequency features (shown in ‘HF PCA’), and dynamic fusion of all features ensures the final aggregated features’ PCA maintains clear semantic edges and successfully filters out artifacts. These visualizations further enhance Earth-Adapters’ explainability.
Component Effectiveness Ablation. In Table 4, ablation studies on MoA’s components and adapter numbers show that using only HF or LF adapters decreases performance compared to spatial adapters, with HF causing a larger drop. This aligns with PCA visualizations in 1 (b) and 3, where HF features have more artifacts and noise, while LF features are smooth with clear global semantics. Thus, using only HF severely damages performance. Although only using LF features outperforms HF features (50.9%mIoU), it still underperforms spatial features (56.9% mIoU) due to lacking semantic details. After that, we also conduct more experiments on the number of Spatial adapters, and reveal the best composition is the one spatial adapter with one HF and LH adapters.
Performance-Speed Trade-Off Ablation. We compare the parameters, speed, and performance between FFT, Rein, and Earth-Adapter in Table 5. Significantly, at the same time of close training and inference speed with Rein, Earth-Adapter achieves better performance (9.0% mIoU improvement), with a smaller parameter scale.
Hyper-Parameter Sensitive Ablation. In this section, we conduct an ablation study to test Earth-Adapter with two hyper-parameter variations on both the P2V (DA) and U2R (DG) benchmarks. Figure 4 presents the studies of the cut-off frequency with the dimension as 16, and Figure 5 illustrates the studies of the projector dimension in the Adapter with the cut-off frequency as 0.3. Both experiments show a consistent trend: Earth-Adapter shows greater robustness in the DA than in the DG setting. We posit that this occurs because the DA task inherently possesses higher certainty compared to DG. In the DA, even when hyper-parameters are altered, Earth-Adapter can effectively learn from the augmented target domain data, thereby maintaining performance stability. Despite DG’s uncertainty, Earth-Adapter shows a relatively stable performance in a certain parameter range. For instance, when the cut-off frequency is between 0.3 and 0.4, performance fluctuations are minimal. Similarly, for the projector dimension, performance is stable across a broad range from 16 to 256 and only fluctuates when the dimension is 512. Importantly, there are no other hyper-parameters requiring tuning. These results underscore Earth-Adapter’s superiority in handling parameter variations, offering a reliable solution for both DA and DG.
Layer Adapter Earth-Adapter
U2R (DA) R2U (DG) U2R (DA) R2U (DG)
Frozen 11.8 57.1 11.8 57.1
[0,1,2,3,4,5] 44.7 57.4 47.6 59.0
[18,19,20,21,22,23] 46.0 58.1 50.0 58.5
Table 7: Layer effect comparison between Adapter and Earth-Adapter. The results show Earth-Adapter learns better under the same layer setting.
Spatial Adapter MoA Frequency Layer Cutoff Frequency U2R (DA) V2P (DA) P2V (DG) R2U (DG)
- - - 48.6 60.1 61.0 58.6
Full 0.3 47.0 60.9 61.3 58.1
shallow 3 0.3 48.3 60.1 63.8 58.0
shallow 6 0.3 47.6 60.1 63.8 59.0
deep 3 0.3 47.4 61.6 63.7 58.7
deep 6 0.3 48.0 60.6 64.9 58.5
deep 3 / deep 6 0.2 / 0.3 50.0 62.2 64.9 59.0
Table 8: Ablation of Frequency Adapter with different activation layers. The results demonstrate that the Earth-Adapter tends to be more beneficial when used in deeper network layers. The results in the gray shadow mean the best performance using deep 3 or 6 layers and 0.2 or 0.3 cutoff frequencies.
4.4 Analysis of Overall Parameter Configuration.
The detailed hyperparameter configurations for each benchmark are presented in Table 6. Overall, our method demonstrates exceptional robustness, exhibiting relatively low sensitivity to parameter variations. Below we provide an in-depth analysis of the hyperparameter effects.
Effect of Adapter dimension. The Earth-Adapter achieves optimal performance across varying dimensions on different benchmarks: P2V (DA) attains peak performance (67.3% mIoU at dim=64), P2V (DG) reaches its highest score (64.9% at dim=24), V2P (DA) achieves 62.2% mIoU at dim=32, while V2P (DG) attains its best performance at dim=64. A notable pattern emerges where domain generalization (DG) tasks consistently require larger-dimension adapters—V2P (DG) and R2U (DG) both peak at dim=64—whereas domain adaptation (DA) tasks achieve optimal results with smaller dimensions (V2P (DA) at dim=32 and R2U (DA) at dim=16). We attribute this phenomenon to DG tasks demanding greater parameter capacity to facilitate the foundation model’s extraction of more comprehensive semantic representations, thereby significantly boosting cross-domain generalization performance.
Effect of cutoff frequency. The cutoff frequency in the frequency domain serves as the boundary between high and low frequencies. Empirical evidence suggests that a relatively low cutoff value is generally preferable. In our experiments, we evaluated performance at cutoff frequencies of 0.2 and 0.3. Different benchmarks achieved optimal results at different frequencies: six benchmarks (P2V (DA), P2V (DG), V2P (DG), R2U (DA), R2U (DG), and U2R (DG)) attained peak performance at a cutoff ratio of 0.3, while V2P (DA) and U2R (DA) performed best at a ratio of 0.2.
Effect of Frequency Adapter layer. We also conducted experiments on the configuration of frequency-domain adapters across different layers. Our study primarily focused on the Dinov2-Large model, a 24-layer Vision Transformer architecture. We evaluated four distinct adapter configurations: {0,1,2}, {0,1,2,3,4,5}, {21,22,23}, and {18,19,20,21,22,23}. As shown in Table 6, different benchmarks achieved optimal performance with varying frequency-domain adapter configuration. For instance, P2V (DA) attained its peak performance of 67.7% mIoU with adapters in 0,1,2, while P2V (DG) achieved its best results with the {18,19,20,21,22,23} configuration. Similar to the adapter dimension findings, the frequency adapter layers demonstrate task-dependent optimization patterns. Notably, three DG tasks (P2V (DG), V2P (DG), and U2R (DG)) consistently performed best with adapters in the layers {18,19,20,21,22,23}, suggesting that domain generalization (DG) requires more adapter parameters to capture general semantic representations to improve cross-domain performance.
4.5 Discussion on Adapter Layer
Layer-wise comparison between Adapter and Earth-Adapter. We analyze the layer-wise behavior of Adapter and compare the performance of Adapter and Earth-Adapter under different layer configurations (Table 7). While prior work [71] suggests that deeper layers capture more semantically meaningful features, our Earth-Adapter demonstrates superior performance in both shallow-layer (first six) and deep-layer (last six) fine-tuning. For instance, in the U2R (DA) task, Earth-Adapter (shallow fine-tuning) outperforms the frozen model by 35.8% mIoU, while in R2U (DG), it surpasses the frozen baseline by 1.9% mIoU. Similar improvements are observed for deep-layer fine-tuning, achieving +38.2% mIoU (U2R (DA)) and +1.6% mIoU (U2R (DG)). Moreover, compared to Adapter, Earth-Adapter consistently shows better adaptation—+2.9% mIoU (U2R (DA), shallow) and +0.4% mIoU (R2U (DG), deep). These results highlight Earth-Adapter’s robustness and generalization capability across different layer-wise configurations.
Detailed analysis of parameter configuration of Frequency Adapter. We conduct a comprehensive analysis of the impact of frequency-domain layers on model performance, as shown in Table 8. In our default setting, the cutoff frequency of Earth-Adapter is set to 0.3. Our experiments reveal that these frequency layers play a crucial role in the overall performance of Earth-Adapter. Notably, integrating the Frequency Adapter results in a slight performance drop in two benchmarks: U2R (DA) ($-1.6\%$ mIoU) and R2U (DG) ($-0.5\%$ mIoU). However, with an efficient parameter search, Earth-Adapter still surpasses the Spatial Adapter. As shown in the table, Earth-Adapter achieves 50.0% mIoU in the U2R (DA) task, outperforming Spatial-Adapter by 1.4% mIoU. Similarly, it attains 59.0% mIoU in the R2U (DG) task, exceeding Spatial Adapter by 0.4% mIoU. A key observation is that Earth-Adapter exhibits sensitivity to frequency layer configurations and cutoff frequencies under certain benchmarks—an aspect we aim to refine in future work.
5 Conclusion
In this paper, we first highlight the performance drop when transferring existing PEFT methods to the RS cross-domain experiments. Detailedly, we analyze these problems through the research [12] and our artifacts-related findings, then raise that existing methods can not effectively handle the artifacts appearing in RS images. To tackle this, we introduce Earth-Adapter, our RS-specialized PEFT method. It is a simple-yet-effective and the first PEFT method for cross-domain RS semantic segmentation. In the experiment section, we first comprehensively evaluate Earth-Adapter’s performance and compare with other methods on widely-used benchmarks. Secondly, we analyze Earth-Adapter’s effectiveness with different backbones, as well as the trade-off between parameters, speed, and performance. Finally, we visualize the decoder’s predictions and expert outputs. Our results demonstrate Earth-Adapter’s power, efficiency, robustness, and explainability.
However, there is still space for improvement. For example, it’s still unclear on which backbone layers it should be applied, and how to further enhance the utilization of frequency features. In our subsequent work, we’ll pursue more potential optimization. The code will be public soon, and we hope Earth-Adapter will be helpful to this community.

### v4

@@
-with Mixture of Frequency Adaptation
+with a Frequency-Guided Mixture of Adapters
 Abstract
-Parameter-Efficient Fine-Tuning (PEFT) is a technique that allows us to adapt powerful Foundation Models (FMs) to diverse downstream tasks while preserving and unleashing their inherent capabilities.
-However, we have observed that existing PEFT methods, which are often designed with natural imagery in mind, struggle when applied to Remote Sensing (RS) scenarios.
-This is primarily due to their inability to handle artifact influences, a problem particularly severe in RS image features.
-To tackle this challenge, we introduce Earth-Adapter, the first PEFT method specifically designed for RS artifacts conquering.
-Earth-Adapter introduces a novel Mixture of Frequency Adaptation process that combines a Mixture of Adapter (MoA) with Discrete Fourier Transformation (DFT).
-By utilizing DFT, Earth-Adapter can decompose features into different frequency components, precisely separating artifacts from original features.
-The MoA then dynamically assigns weights to each adapter expert, allowing for the combination of features across various frequency domains.
-These simple-yet-effective approaches enable Earth-Adapter to more efficiently overcome the disturbances caused by artifacts than previous PEFT methods, significantly enhancing the FMs’ performance on RS scenarios.
-Experiments on Domain Adaptation (DA), and Domain Generalization (DG) semantic segmentation benchmarks showcase the Earth-Adapter’s effectiveness.
-Compared with baseline Rein, Earth-Adapter significantly improves 9.0% mIoU in DA and 3.1% mIoU in DG benchmarks.
-Our code will be released at Earth-Adapter.
-1 Introduction
-Tremendous last year breakthroughs in Foundation Models (FMs) [14, 45, 54, 40, 25, 55, 22, 3] have ignited transformative progress in Computer Vision (CV), demonstrating unprecedented capabilities across both high-level and low-level vision tasks.
-Based on these FMs, considerable Remote Sensing (RS) research groups have also constructed large-scale, state-of-the-art (SOTA), and impressive Geospatial Foundation Models (GFMs) [5, 51, 21, 65, 70, 43, 36, 23, 50, 38], significantly promoting the development of RS imagery interpretation field.
-However, when adapting these FMs to downstream tasks, how to efficiently keep and unleash the power of FMs becomes a key problem.
-Under this context, Parameter-Efficient Fine-Tuning (PEFT) methods [31, 37, 28, 8, 20] become the pivotal method due to their superior parameter-performance trade-off.
-Existing PEFT methods designed for language or nature imagery have made great progress with excellent works, such as LoRA [31], Visual Prompt Tuning (VPT) [37].
-However, most of them show performance drops in RS cross-domain experiments.
+Vision Foundation Models (VFMs), while powerful, often struggle in Remote Sensing (RS) segmentation tasks when combined with existing Parameter-Efficient Fine-Tuning (PEFT) methods.
+We observe that this limitation primarily arises from their inability to effectively handle the pervasive artifacts in RS imagery.
+To address this, we introduce Earth-Adapter, the first PEFT method specifically designed for RS artifact mitigation.
+Earth-Adapter introduces a novel Frequency-Guided Mixture of Adapters (MoA) approach, structured around a “divide and conquer” strategy.
+It first utilizes Discrete Fourier Transformation (DFT) to ”divide” features into distinct frequency components, thereby effectively isolating artifact-related information from semantic signals.
+Subsequently, to “conquer” these artifacts, MoA independently optimizes features within different subspaces and dynamically assigns weights via a router to aggregate the refined representations.
+This enables adaptive refinement of the VFM’s representation space to mitigate the impact of artifacts.
+This simple yet highly effective PEFT method demonstrably mitigates artifacts and significantly enhances VFMs’ performance on RS segmentation tasks.
+Extensive experiments demonstrate Earth-Adapter’s effectiveness on in-domain semantic segmentation (SS), as well as Domain Adaptive (DA) and Domain Generalized (DG) semantic segmentation tasks.
+Compared with the baseline Rein, Earth-Adapter significantly improves mIoU by 1.2% in SS, 9.0% in DA, and 3.1% in DG benchmarks.
+Our code is available at https://github.com/VisionXLab/Earth-Adapter.
+Introduction
+Figure 1: Performance across various remote sensing image segmentation benchmarks between Frozen VFM (DINOv2-L), Rein (Baseline) and the proposed Earth-Adapter.
+Vision Foundation Models (VFMs) have made significant strides in recent years, driven by both natural language supervision paradigms (e.g., CLIP (Radford et al.
+2021), ALIGN (Jia et al.
+2021), SigLIP (Zhai et al.
+2023)) and self-supervised paradigms (e.g., MAE (He et al.
+2022a), DINOv2 (Oquab et al.
+2023)).
+Pre-training on ultra-large-scale datasets empowers VFMs with robust zero-shot comprehension capabilities, leading to their extensive adoption across diverse tasks.
+These include classical computer vision tasks (classification (Zhou et al.
+2022), detection (Zareian et al.
+2021; Minderer et al.
+2022), segmentation (Chen et al.
+2024a)) as well as multimodal learning (Liu et al.
+2023).
+Simultaneously, the conventional pre-train-then-fine-tune paradigm is evolving.
+When adapting VFMs to downstream tasks, the key challenge becomes how to efficiently preserve and unleash their inherent capabilities.
+Under this context, Parameter-Efficient Fine-Tuning (PEFT) methods (Hu et al.
+2021; Jia et al.
+2022; Houlsby et al.
+2019a; Gong et al.
+2024) emerge as the pivotal solution due to their superior parameter-performance trade-off.
+Existing PEFT methods designed for language or nature imagery have made great progress with excellent works, such as LoRA (Hu et al.
+2021), Visual Prompt Tuning (VPT) (Jia et al.
+2022).
+However, most exhibit significant performance degradation in remote sensing (RS) segmentation tasks when integrated with DINOv2 (Oquab et al.
+2023) (experimental evidence provided in later sections).
 We argue the main reason lies in the influences of artifacts in the features.
 As shown in Figure 2 (a), we show the PCA visualization of features of DINOv2-L, which contains obvious redundant artifacts.
-Although a part of existing PEFT methods also are shown to be capable of handling artifacts [12], they still underestimate the artifact disturbances in RS imagery.
-The difference in the distribution of artifacts between natural and RS images is significant.
-In natural images, artifacts typically surround foreground objects [12], such as humans or animals, and the disturbances they cause are relatively limited.
+We have also observed that artifacts in natural images exhibit distinct characteristics compared to those in RS images.
+In natural images, artifacts typically surround foreground objects (Darcet et al.
+2024), such as humans or animals, and the disturbances they cause are relatively limited.
 In contrast, RS images, due to their overhead perspective, lack centralized subjects and contain multiple coexisting multi-scale targets.
 For example, a single RS image may simultaneously include large-scale agricultural regions and fragmented road networks.
-As a result, artifacts are almost situated everywhere in RS images shown in Figure 1 (a), causing severe interference to domain-invariant representation learning, which is pivotal for cross-domain scenarios.
+As a result, artifacts are almost situated everywhere in RS images shown in Figure 2 (a), causing severe interference to pixel-level feature extraction, which is pivotal for segmentation tasks.
 […]
-In our view, to address the influence of artifacts, one of the core ideas is to ‘divide-and-conquer’.
-Drawing on the principle that high-frequency (HF) signals encode local details (e.g., edge, contour) while low-frequency (LF) signals encode global structure (e.g., object surface), we leverage Discrete Fourier Transformation (DFT) as the ‘divide’ step.
-As illustrated in Figure 1 (b), we successfully leverage the DFT effectively to isolate artifacts into HF signals and splits coherent global features into LF signals.
-Subsequently, to ‘Conquer’ artifact-related issues, we adopt a Mixture of Adapter (MoA) architecture for PEFT design.
-The MoA router dynamically assigns weights to adapter experts, selecting the most relevant features for processing in each iteration.
-In this way, all features are aggregated into optimized representations, mitigating artifacts disturbances.
-Based on our motivation and findings, we further present Earth-Adapter, a simple-yet-effective RS-specialized PEFT method designed for RS artifact conquering.
-Specifically, Earth-Adapter consists of a novel Mixture of Frequency Adapter architecture consisting of Mixture of Adapter (MoA) with Discrete Fourier Transformation (DFT).
-Earth-Adapter first disentangles geospatial features by DFT, and leverages multiple experts to conquer artifacts by adaptively aggregating and learning features with suitable frequency.
-In this fashion, Earth-Adapter can effectively alleviate the artifacts’ disturbances for robust domain-invariant feature learning.
-Then gradually introduces FMs to cross the geospatial gaps from the frequency perspective.
-In experiments, we meticulously test Earth-Adapter on Domain Adaptation (DA), and Domain Generalization (DG) semantic segmentation tasks across five widely-used datasets.
-Comparing with existing PEFT methods, Earth-Adapter achieves State-Of-The-Art (SOTA) performance on DA and DG benchmarks.
-In DA and DG benchmarks, Earth-Adapter outperforms our baseline Rein with 8.9% and 3.2% mIoU, respectively.
-Additionally, we also conduct in-depth analysis of the potential of Earth-Adapter.
+To address artifacts and improve VFM performance in RS segmentation tasks, we propose a “divide-and-conquer” strategy.
+In the “divide” stage, based on the observation that high-frequency (HF) signals capture local details while low-frequency (LF) signals encode global structures, we apply Discrete Fourier Transformation (DFT) to separate HF and LF components, isolating artifacts in the HF domain.
+In the “conquer” stage, we design the Mixture of Adapters (MoA), which adjusts features in their respective frequency domains and then uses a router to dynamically assign weights for aggregating the corrected features.
+These aggregated features are further combined with the original frozen VFM features through skip connections and dynamic scaling coefficients, preserving the strong feature extraction ability of the underlying model.
+This unique pipeline forms Earth-Adapter, a PEFT method tailored for RS semantic segmentation.
+We comprehensively evaluate Earth-Adapter on 12 established benchmarks across three settings—in-domain (SS), domain adaptive (DA), and domain generalized (DG)—and demonstrate its effectiveness.
+Compared with existing PEFT methods, Earth-Adapter achieves state-of-the-art (SOTA) performance on SS, DA and DG benchmarks and outperforming our baseline Rein (Wei et al.
+2024) by 1.2% , 9.0% and 3.1% mIoU, respectively.
+We also conduct an in-depth analysis of Earth-Adapter’s design and potential.
 The core contributions can be summarized as follows:
-  • We propose a novel perspective to alleviate artifact-disturbance problems in RS imagery by frequency-guided divide-and-conquer strategy.
-  • We introduce a simple-yet-effective PEFT method, Earth-Adapter. It is also the first RS-specialized PEFT approach designed for cross-domain semantic segmentation tasks.
-  • Extensive comparisons validate the SOTA performance of Earth-Adapter on both DA and DG benchmarks.
-More ablation studies also demonstrate the high efficiency, robustness, and explainability of Earth-Adapter.
-2 Related Work
-2.1 RS Cross-Domain Semantic Segmentation
-As a classical and fundamental task in the realm of computer vision, semantic segmentation [46, 57] not merely holds profound practical significance, but also provides an interface for researchers to dissect the high-level semantic representation of images.
-With the growing need for applying models to out-of-domain data, cross-domain segmentation, including domain adaptation (DA) [29, 30, 19, 41, 33] and domain generalization (DG) [67, 21], has gained prominence.
-DA and DG are the two main paradigms, where DA aims to adapt a model from a labeled source domain to an unlabeled target domain, while DG focuses on training a model on only the source domain to generalize to unseen target domains.
-In the RS field, cross-domain task is crucial due to the diversity of data, and lots of DA methods have been applied to improve the generalization ability of segmentation models, with applications in urban planning [56], land resource management [60], and environmental protection [58].
-Compared with DA, the number of DG research [34, 42] in RS is much less than DA.
-Although these works [72, 35, 16] make contributions, most of them only focus on specialized models rather than leveraging the power of FMs, leading to limited cross-domain capabilities.
-2.2 Geospatial Foundation Models
-The emergence of Large Language Models (LLMs) [4] has prompted a shift in computer vision, leading to the development of Multimodal Large Language Models (MLLMs) [48, 54, 44, 9] and Vision Foundation Models (VFMs) [45, 52, 24, 40].
-FMs [3] in RS called GFMs [64, 26, 63, 62, 6] also can be divided into MLLMs and VFMs.
-Among MLLMs, RemoteCLIP [43] is the first exploration, supporting classification, retrieval, and object counting through a contrastive learning framework based on CLIP [54].
-MetaEarth [70] advances diffusion models to generate multi-resolution RS images for any specified region based on both image and text conditions.
-Skysense [23] captures spatiotemporal information and performs well across diverse scenarios.
-Although these FMs obtain strong power by large-scale pertaining, how to effectively unleash their power to diverse downstream tasks is still a challenge.
-2.3 PEFT in RS
-PEFT approaches, such as adapter [28, 27, 31] are proposed to fine-tune large pre-trained models to downstream tasks in Natural Language Processing (NLP) by introducing light-weight learnable parameters.
-Subsequently, the PEFT has begun to be transferred to vision[69, 1, 39] tasks and rapidly become the focus of researchers.
-Visual Prompt Tuning (VPT) [37] first introduced ‘prompt tuning’ into the vision area as learnable vectors.
-The thought of this paradigm is also similar to the adversarial re-programming [15].
-Currently, cross-domain works also focus on PEFT, such as [18] which is the first prompt tuning method in Domain Adaptation (DA) [59, 17, 20] and [67] is the first work leveraging PEFT to fine-tuning VFMs for Domain Generalization (DG) [2].
+  • We observe that VFMs are prone to artifact interference in the high-dimensional feature space of RS imagery, which hinders their adaptation to RS segmentation tasks.
+  • To address this, we propose Earth-Adapter, the first PEFT approach tailored specifically to mitigate artifact-related issues.
+  • We develop a Frequency-Guided Mixture of Adapters (MoA) that effectively isolates artifacts within distinct frequency subspaces, coupled with a dynamic router that adaptively fuses adapter outputs to optimize pixel-level VFM features for RS imagery.
+  • Extensive experiments across diverse RS segmentation benchmarks demonstrate the effectiveness of Earth-Adapter.
+Related Work
+RS Semantic Segmentation
+Semantic segmentation (Long, Shelhamer, and Darrell 2015; Shelhamer, Long, and Darrell 2016; Chen et al.
+2017) aims to classify every pixel in an image.
+RS segmentation is a critical task in numerous real-world applications, including land cover mapping, urban planning, and environmental monitoring (He et al.
+2022b).
+Cross-domain task is crucial due to the diversity of data, and lots of DA methods have been applied to improve the generalization ability of segmentation models(Rui and Jintao 2020; Tong et al.
+2020).
+Compared with DA, the number of DG research (Iizuka, Xia, and Yokoya 2023; Liang et al.
+2024) in RS is much less than DA.
+Although these works (Zhu et al.
+2023) make contributions, most of them only focus on specialized models rather than leveraging the power of VFMs, leading to limited cross-domain capabilities.
+Vision Foundation Models
+Vision Foundation Models (VFMs) have made significant progress and have exerted considerable influence across various AI-related fields.
+Among them, vision-language foundation models—such as CLIP (Radford et al.
+2021), ALIGN (Jia et al.
+2021), SigLIP (Zhai et al.
+2023), and InternViT (Chen et al.
+2024b)—are trained on billions of image-text pairs to learn joint visual-textual representations.
+These models have been widely applied to a broad range of vision and multimodal tasks, including classification (Zhou et al.
+2022), detection (Zareian et al.
+2021; Minderer et al.
+2022), segmentation (Chen et al.
+2024a), image-text retrieval (Radford et al.
+2021), image captioning (Dai et al.
+2023), and large multimodal models (LMMs) (Luo et al.
+2024).
+On the other hand, self-supervised models can also learn rich visual representations from large-scale unlabeled data.
+MAE (He et al.
+2022a) learns generalized visual features by reconstructing masked images from partial inputs.
+MoCo (Chen, Xie, and He 2021) employs a momentum encoder and contrastive learning to extract discriminative representations.
+DINOv2 (Oquab et al.
+2023) combines self-distillation with contrastive learning to produce scalable, transferable features for various downstream tasks.
+Beyond these various VFMs, a key challenge remains in effectively unlocking their potential to enhance performance on downstream tasks.
+Earth-Adapter is a pioneering work for RS segmentation tasks
+PEFT in RS
+PEFT approaches, such as adapter (Houlsby et al.
+2019a, b; Hu et al.
+2021) are proposed to fine-tune large pre-trained models to downstream tasks in Natural Language Processing (NLP) by introducing light-weight learnable parameters.
+Subsequently, the PEFT has begun to be transferred to vision(Yin et al.
+2023; Agiza, Neseem, and Reda 2024) tasks and rapidly become the focus of researchers.
+Visual Prompt Tuning (VPT) (Jia et al.
+2022) first introduced ‘prompt tuning’ into the vision area as learnable vectors.
+The thought of this paradigm is also similar to the adversarial re-programming (Elsayed, Goodfellow, and Sohl-Dickstein 2018).
+Currently, cross-domain works also focus on PEFT, such as (Ge et al.
+2023) which is the first prompt tuning method in Domain Adaptation (DA) (Sun et al.
+2023; Gao et al.
+2022; Gong et al.
+2024) and (Wei et al.
+2024) is the first work leveraging PEFT to fine-tuning VFMs for Domain Generalization (DG) (Bi et al.
+2024).
 In the RS field, the exploration of PEFT is also vigorous.
-For example, UPetu [13] addresses storage for dense prediction tasks.
-TEA [32] uses an adapter network and top-down guidance for detection and segmentation tasks.
-[53] also leverage LoRA to solve oriented detection task.
-However, for the cross-domain semantic segmentation tasks, the research of PEFT is relatively blank, and existing methods can not handle the influence of artifacts.
-Earth-Adapter is the first PEFT for this area and can effectively tackle the artifacts.
-3 Method
-3.1 Preliminary
-We first denote the input RS images as $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, the corresponding label of $\mathbf{x}$ as $\mathbf{y}\in\mathbb{R}^{H\times W\times K}$, and the semantic segmentation prediction of $\mathbf{x}$ as $\mathbf{y}^{\prime}\in\mathbb{R}^{H\times W\times K}$, where $K$ represents the number of semantic classes.
+For example, UPetu (Dong, Gu, and Liu 2024) addresses storage for dense prediction tasks.
+TEA (Hu et al.
+2024) uses an adapter network and top-down guidance for detection and segmentation tasks.
+(Pu and Xu 2025) also leverage LoRA to solve oriented detection task.
+However, for the RS semantic segmentation tasks, the research of PEFT is relatively blank, and existing methods can not handle the influence of artifacts.
+Earth-Adapter is the first PEFT designed to tackle artifacts, thereby unlocking the potential of VFMs for remote sensing segmentation tasks.
+Method
+Preliminary
+We first denote the input RS image as $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, the corresponding label of $\mathbf{x}$ as $\mathbf{y}\in\mathbb{R}^{H\times W\times K}$, and the semantic segmentation prediction of $\mathbf{x}$ as $\mathbf{y}^{\prime}\in\mathbb{R}^{H\times W\times K}$, where $K$ represents the number of semantic classes.
 Here, we’ll illustrate our innovative Earth-Adapter overview.
-The proposed Earth-Adapter is composed of two key components: the DFT and MoA.
-MoA consists of a Spatial Adapter, a LF Adapter, a HF Adapter, and a router.
+The proposed Earth-Adapter is composed of two key components: the MoA and Router.
+MoA consists of a Spatial Adapter, a low-frequency (LF) Adapter, a high-frequency (HF) Adapter.
 We employ $R_{\xi}$, parameterized by $\xi$, to denote the router and $E^{i}_{\epsilon},\;i\in\{1,2,3\}$, parameterized by $\epsilon$, to represent the three adapter experts.
-For the network architecture, we adopt the framework introduced by Rein [67], integrating DINOv2 [52] as the backbone, represented as $f_{\phi}$ and parameterized by $\phi$, alongside Mask2Former [10] as the decoder, denoted by $f_{\theta}$ and parameterized by $\theta$.
-Section 3.2 illustrates the Earth-Adapter’s design, while Section 3.3 elaborates on our cross-domain training framework.
-3.2 Details of Earth-Adapter
-The core idea of Earth-Adapter is the seamless integration of highly efficient, learnable layers with minimal parameters into the frozen FMs.
+For the network architecture, we adopt the framework introduced by Rein (Wei et al.
+2024), integrating DINOv2-L (Oquab et al.
+2023) as the backbone, represented as $f_{\phi}$ and parameterized by $\phi$, alongside Mask2Former (Cheng et al.
+2022) as the decoder, denoted by $f_{\theta}$ and parameterized by $\theta$.
+The following presents the details of Earth-Adapter, and we will elaborate on our training framework in the Training Framework section.
+Details of Earth-Adapter
+Earth-Adapter aims to adapt VFMs to RS image semantic segmentation tasks with minimal learnable parameters.
 Optimization objective.
-Before employing the Earth-Adapter, the optimization objective of fine-tuning FMs is to identify a set of parameters that minimize the loss of the entire model on the downstream task:
+Before employing the Earth-Adapter, the optimization objective of fine-tuning VFMs is to identify a set of parameters that minimize the loss of the entire model on the downstream task:
 $\mathop{\arg\min}_{\theta}\,\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\mathcal{L}_{\text{seg}}\left(f_{\theta}(f_{\phi}(\mathbf{x})),\mathbf{y}\right),$ (1)
-where $\mathcal{D}$ is the dataset, $f_{\phi}\cdot f_{\theta}$ is the segmentation model, and $\mathcal{L}_{seg}$ is the loss function, here we use the default CE loss.
-After activating our Earth-Adapter, the optimization objective becomes:
+where $\mathcal{D}$ is the dataset and $\mathcal{L}_{seg}$ is the loss function, here we use the default CE loss. After incorporating our Earth-Adapter, the optimization objective becomes:
 $\mathop{\arg\min}_{\theta,\epsilon,\xi}\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\mathcal{L}_{\text{seg}}\left(f_{\theta}\left(\underbrace{R_{\xi}\circ E_{\epsilon}}_{\text{Earth-Adapter}}(f_{\phi}^{*}(\mathbf{x}))\right),\mathbf{y}\right).$ (2)
-During training, the original parameters of the backbone are kept frozen (denoted as $f_{\phi}^{*}$), ensuring they remain unaltered.
+During training, the original parameters of the backbone (VFM) are kept frozen (denoted as $f_{\phi}^{*}$).
 In the following description, we denote the segmentation network combined with Earth-Adapter as $G_{\theta,\xi,\epsilon}$.
-Parallel Frequency Adaptation.
-Let $\mathbf{F}_{i}\in\mathbb{R}^{n\times c}$ denote the visual feature at the $i$-th layer of the backbone, where $n$ represents the token sequence length and $c$ denotes the channel dimension.
+Mixture of Adapters.
+Let $\mathbf{F}_{i}\in\mathbb{R}^{(hw)\times c}$ denote the visual feature at the $i$-th layer of the backbone, where $hw$ represents the token sequence length and $c$ denotes the token dimension.
 Our frequency adaptation operates through parallel processing streams: The first Spatial Adapter employs a low-rank projection to refine the spatial feature:
 $\Delta\mathbf{F}^{spatial}_{i}=\text{Adapter}_{1}^{i}(\mathbf{F}_{i}^{T}),$ (3)
-where $\Delta\mathbf{F}^{spatial}_{i}$ means the processed i-th layer Spatial-Domain features, and $\textbf{Adapter}_{1}$ represents a nonlinear mapping layer composed of two low-rank matrices:
+where $\Delta\mathbf{F}^{spatial}_{i}$ means the processed i-th layer Spatial-Domain features, and $\textbf{Adapter}_{1}$ represents a nonlinear mapping layer composed of two low-rank matrices and an activation:
 $\textbf{Adapter}_{i}=\textbf{W}_{up}(\text{Relu}(\textbf{W}_{down}(\cdot))).$ (4)
-The Frequency Adapter decompose features through spectral analysis.
-We first reshape $\mathbf{F}^{spatial}$ to $({C,H,W})$ and apply DFT on spatial features as $\mathcal{FT}(\mathbf{F}_{spatial})$.
+The Frequency Adapters consist of a HF Adapter (high-frequency subspace) and a LF Adapter (low-frequency subspace), which fine-tune features in specific frequency subspaces derived from a 2D Discrete Fourier Transform (DFT) decomposition.
+We first reshape $\mathbf{F}^{spatial}$ to (C,H,W)({C,H,W}) and apply DFT on spatial features as $\mathcal{FT}(\mathbf{F}_{spatial})$.
 When splitting the frequency domains, we employ a fixed frequency cutoff $\rho$ to decompose the spectrum into high and low frequency components.
@@
 $\Delta\mathbf{F}^{high}_{i}=\text{Adapter}_{3}^{i}(\mathbf{F}^{high}_{i}).$ (9)
-Mixture of Adapters.
-We implement dynamic feature fusion through a Router that learns optimal combinations of feature adjustment according to the original visual features.
+The Frequency Adapters shares the same structure as the Spatial Adapter, as defined in Equation 4.
+Dynamic Router.
+We implement dynamic feature aggregation through a Router that learns optimal combinations of feature adjustment according to the original visual features.
 The Router weight is computed by channel-wise attention:
@@
 where $\mathbf{w}_{i}$ represents weights for spatial, LF, and HF components. The final feature adjustment is calculated as:
-$\Delta\mathbf{F}_{i}=\alpha_{i}\sum_{k=1}^{3}\mathbf{w}^{(k)}\Delta\mathbf{F}^{(k)}_{i},$ (11)
-where $\alpha_{i}$ is a learnable scaling parameter with small initial value, and $k\in\{spatial,low,high\}$. The complete feature optimization also maintains residual structure:
+$\Delta\mathbf{F}_{i}=\alpha_{i}\sum_{k=1}^{3}\mathbf{w}_{i}^{(k)}\Delta\mathbf{F}^{(k)}_{i}$, (11)
+where $\alpha_{i}$ is a learnable scaling parameter with small initial value, and $k\in\{spatial,low,high\}$. The frozen features and the refined features are fused via a skip connection:
 $\bar{\mathbf{F}_{i}}=\mathbf{F}_{i}+\Delta\mathbf{F}_{i}.$ (12)
-3.3 Training FrameWork
-In this study, we explore the application of Earth-Adapter across two distinct tasks: DG and UDA Semantic Segmentation.
-For DG Semantic Segmentation tasks, we train Earth-Adapter with Mask2former in end-to-end supervised learning as Equation 2.
-For UDA Semantic Segmentation, we use the self-training framework introduced in DACS [61] without using target domain labels.
-The pseudo labels for target domain images are produced by a teacher network $G_{\theta,\xi,\epsilon}^{\dagger}$:
-$\mathbf{\bar{y}_{T}}^{(j,k)}=[k=\mathop{\arg\max}_{k^{\prime}}G_{\theta,\xi,\epsilon}^{\dagger}(\mathbf{x_{T}})))^{(j,k^{\prime})}],$ (13)
-where []delimited-[][$\cdot$][ ⋅ ] denotes the Iverson bracket, and $j\in HW$ is the spatial index of the pixel in the image.
-The teacher network remains inactive during the training process, with its parameters being updated based on an exponential moving average (EMA) of the student model’s parameters:
-$G_{\theta,\xi,\epsilon,t+1}^{\dagger}=\alpha G_{\theta,\xi,\epsilon,t}^{\dagger}+(1-\alpha)G_{\theta,\xi,\epsilon,t}$ (14)
-Here $t$ denotes the $t_{th}$ training iteration.
-The hyperparameter $\alpha$ controls the update speed of the teacher network’s parameters.
-After obtaining the pseudo-labels, we perform category-wise mixing of samples from the source and target domains:
-$\mathbf{x}_{mix}$ $=\mathbf{x_{S}}\odot\mathcal{M}+\mathbf{x_{T}}\odot(1-\mathcal{M})$ (15)
-$\mathbf{y}_{mix}$ $=\mathbf{y_{S}}\odot\mathcal{M}+\mathbf{\bar{y}_{T}}\odot(1-\mathcal{M})$
-where $(\mathbf{x_{S}},\mathbf{y_{S}})$ is the data sample from the source domain, and $\mathcal{M}$ is generated by randomly selecting half of the categories based on the labels of the source domain $\mathbf{y_{S}}$.
-The Optimization objective for the UDA task if defined by:
+$\bar{\mathbf{F}_{i}}$ is then forwarded to the subsequent Transformer block to continue the layer-wise processing.
+Training Framework
+In this work, we explore the application of Earth-Adapter across three distinct semantic segmentation subtasks: semantic segmentation (SS), domain adaptive (DA) semantic segmentation and domain generalized (DG) semantic segmentation .
+For SS and DGSS, we train Earth-Adapter with Mask2former (Cheng et al.
+2022) in end-to-end supervised learning as described in Equation 2.
+For DA semantic segmentation, we use the self-training framework introduced in DACS (Tranheden et al.
+2021) without using target domain labels.
+DACS employs an EMA teacher to generate pseudo labels for target-domain images, mixes target and source samples to obtain the mixed domain $(\mathbf{x}_{mix},\mathbf{y}_{mix})$ and then jointly trains on both the source and mixed domains.
+The optimization objective for the DASS if defined by:
 $\mathop{\arg\min}_{\theta,\epsilon,\xi}$ $\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}\cup\mathcal{D}_{mix}}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}),\mathbf{y}\right)$ (13)
-$\quad+\lambda_{uda}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}_{mix}),\mathbf{y}_{mix}\right).$
-$\lambda_{uda}$ is the weight parameter for UDA loss.
-4 Experiments
-Backbone Methods Cross-Domain Adaptation Avg. Cross-Domain Generalization Avg.
+$\quad+\lambda_{da}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}_{mix}),\mathbf{y}_{mix}\right).$
+$\lambda_{da}$ is the weight parameter for DA loss.
+Experiments
+Methods DASS Avg. DGSS Avg.
 P2V V2P R2U U2R P2V V2P R2U U2R
-MiT-B5 [68] DAFormer [29] 64.4 54.8 52.7 42.5 53.6 42.4 41.4 54.2 39.9 44.5
-HRDA [30] 67.6 58.6 53.2 35.3 53.7 33.1 31.1 54.2 39.8 39.6
-DINOv2-L [52] Frozen 21.0 7.2 21.5 11.8 15.4 57.9 49.4 57.1 42.7 51.8
+Previous SOTA methods
+DAFormer 64.4 54.8 52.7 42.5 53.6 42.4 41.4 54.2 39.9 44.5
+HRDA 67.6 58.6 53.2 35.3 53.7 33.1 31.1 54.2 39.8 39.6
+PEFT-based methods with VFM
+Frozen 21.0 7.2 21.5 11.8 15.4 57.9 49.4 57.1 42.7 51.8
+Frozen (with register) 11.3 28.2 36.4 16.1 23.0 59.0 51.1 58.9 44.9 53.5
 Full Fine-Tune 11.3 16.5 23.1 10.6 15.4 12.6 19.6 33.1 19.7 21.3
-Adapter [28] 66.4 59.3 55.9 46.2 57.0 56.0 47.4 57.6 41.8 50.7
-LoRA [31] 17.8 18.8 24.5 26.0 15.7 20.3 25.6 29.3 21.9 24.3
-VPT [37] 66.2 59.3 55.3 48.0 57.2 59.2 52.3 57.4 44.9 53.5
-AdaptFormer [7] 12.9 15.3 25.1 22.2 18.9 14.0 19.3 31.0 15.8 20.0
-Rein (Baseline) [67] 60.2 60.9 52.8 26.0 50.0 60.8 52.5 55.8 43.4 53.1
-Earth-Adapter (Ours) 67.7 (+7.5) 62.2 (+1.3) 55.9 (+3.1) 50.0 (+24.0) 59.0 (+9.0) 64.9 (+4.1) 55.1 (+2.6) 59.0 (+3.2) 45.7 (+2.3) 56.2 (+3.1)
-Improvements over the second best +0.1 +1.3 +0.0 +2.0 +1.8 +4.1 +2.5 +1.4 +0.8 +2.7
-Table 1: Performance Comparison between representative DA models, PEFT methods, and our Earth-Adapter on DA and DG benchmarks.
-The Bold font in the table means the SOTA performance.
-The ‘+’ in the bracket means the improvement compared with our baseline Rein.
-The last line of data means the improvement compared with the second-best performance of all other methods.
-Methods Setting Surf Bldg Vegt Tree Car Clut mIoU (%) Setting Bkgd Bldg Road Water Barr Frst Agri mIoU (%)
-Frozen P2V (DA) 38.6 24.1 11.9 47.2 2.5 1.5 21.0 U2R (DA) 42.6 17.2 6.7 9.6 0.7 5.7 0.0 11.8
-Rein [67] 75.4 88.6 50.3 73.4 59.0 14.5 60.2 27.9 41.9 29.3 51.2 0.8 53.2 0.0 29.2
-Earth-Adapter (Ours) 80.3 89.5 66.3 74.4 61.9 33.6 67.7 56.3 63.4 48.7 66.6 13.7 38.4 63.2 50.0
-Frozen P2V (DG) 70.6 79.9 46.1 72.1 58.6 20.4 57.9 U2R (DG) 56.7 54.8 41.3 67.0 14.6 15.1 49.2 42.7
-Rein [67] 74.4 85.1 51.0 23.0 57.1 24.5 60.8 53.0 57.8 41.2 61.1 9.4 23.4 57.8 43.4
-Earth-Adapter (Ours) 76.3 85.4 59.0 74.1 54.2 40.6 64.9 52.6 55.7 44.1 64.6 17.0 29.6 56.1 46.7
-Table 2: Performance Comparison of Per-Class between Frozen DINOv2-L, Rein, and Earth-Adapter on Potsdam RGB-to-Vaihingen.
-Both Rein and ours are based on the Frozen DINOv2-L.
-The Bold font means the best performance.
+Adapter 66.4 59.3 55.9 46.2 57.0 56.0 47.4 57.6 41.8 50.7
+LoRA 17.8 18.8 24.5 26.0 15.7 20.3 25.6 29.3 21.9 24.3
+VPT 66.2 59.3 55.3 48.0 57.2 59.2 52.3 57.4 44.9 53.5
+AdaptFormer 12.9 15.3 25.1 22.2 18.9 14.0 19.3 31.0 15.8 20.0
+Rein (baseline) 60.2 60.9 52.8 26.0 50.0 60.8 52.5 55.8 43.4 53.1
+Earth-Adapter (Ours) 67.7 62.2 55.9 50.0 59.0 64.9 55.1 59.0 45.7 56.2
+$\Delta$ over baseline method +7.5 +1.3 +3.1 +24.0 +9.0 +4.1 +2.6 +3.2 +2.3 +3.1
+Table 1: Performance (mIoU%) comparison between previous SOTA methods and PEFT-based methods on DA and DG benchmarks.
+Bold indicates the best performance.
+Underlined results denote the second-best.
+The last row shows improvements over the baseline.
+Method P V L i Avg.
+Traditional Methods
+DeepLabV3+ (R-101) 74.8 69.7 51.4 53.4 62.3
+PSPNet (R-101) 74.7 69.2 48.5 57.5 62.5
+UperNet (Swin-B) 75.5 69.4 55.6 67.9 67.1
+Segformer (MiT-B5) 75.3 68.4 55.5 66.8 66.5
+Mask2Former (Swin-B) 75.7 71.2 54.7 64.0 66.4
+PEFT-based methods with VFM
+Frozen 75.6 70.4 54.2 67.3 66.9
+Frozen (with register) 75.7 70.8 56.0 67.5 67.5
+Full Fine-Tune 70.5 64.6 43.9 12.0 47.7
+LoRA 71.9 64.6 51.0 48.5 59.0
+VPT 75.9 70.6 54.6 68.5 67.4
+AdaptFormer 65.4 56.0 42.0 7.5 42.8
+Rein 76.2 70.8 54.9 68.4 67.6
+Earth-Adapter (Ours) 76.7 71.7 56.9 69.8 68.8
+$\Delta$ over baseline method +0.5 +0.9 +2.0 +1.4 +1.2
+Table 2: Performance (mIoU%) comparison between previous SOTA methods and PEFT-based methods on $S$ benchmarks.
+Bold indicates the best performance.
+Underlined results denote the second-best.
+The last row shows improvements over the baseline.
+General VFMs
 Backbone Methods Params P2V(DG) V2P(DG) Avg.
-CLIP [54] (ViT-Large) Frozen 0.0M 43.7 44.4 44.1
-Full Fine-Tune 304.2M 11.4 20.3 15.9
-Earth-Adapter 2.4M 49.7 50.2 50.0 (+5.9)
-DINOv2 [52] (Small) Frozen 0.0M 44.3 43.3 43.8
-Full Fine-Tune 22.1M 41.9 41.8 41.9
-Earth-Adapter 2.4M 47.5 44.8 46.2 (+2.4)
-DINOv2 [52] (Base) Frozen 0.0M 51.4 49.5 50.5
-Full Fine-Tune 86.6M 46.7 49.8 49.8
-Earth-Adapter 2.4M 53.0 51.8 52.4 (+1.9)
-DINOv2 [52] (Large) Frozen 0.0M 57.9 49.4 53.7
-Full Fine-Tune 304.2M 12.6 19.6 16.1
-Earth-Adapter 2.4M 64.9 55.1 60.0 (+6.3)
-Table 3: Ablation studies of Earth-Adapter on different foundation backbones. Both P2V and V2P are the DG benchmarks. Results validate that Earth-Adapter’s effectiveness on diverse backbone networks.
+DINOv2-S
+Frozen 0.0M 44.3 43.3 43.8
+Earth-Adapter 2.6M~9.6M 47.5 44.8 46.2
+DINOv2-B
+Frozen 0.0M 51.4 49.5 50.5
+Earth-Adapter 2.6M~9.6M 53.0 51.8 52.4
+DINOv2-L
+Frozen 0.0M 57.9 49.4 53.7
+Earth-Adapter 2.6M~9.6M 64.9 55.1 60.0
+RS-pretrained VFMs
+Backbone Methods Params P2V(DA) V2P(DA) Avg.
+MTP-L
+Frozen 0.0M 32.0 33.0 32.5
+Earth-Adapter 2.6M~9.6M 40.8 41.0 40.9
+ScaleMAE-L
+Frozen 0.0M 20.0 19.8 19.9
+Earth-Adapter 2.6M~9.6M 20.2 29.6 24.9
+DOFA-L
+Frozen 0.0M 9.9 13.7 11.8
+Earth-Adapter 2.6M~9.6M 26.2 33.2 20.7
+Table 3: Ablation studies of Earth-Adapter on different VFMs. Results validate that Earth-Adapter’s effectiveness on diverse backbone networks.
+Figure 3: Visualization of Predicted Segmentation Maps We Compare Earth-Adapter with the Frozen DINOv2-L backbone and our baseline Rein on eight cross-domain benchmarks.
+For the Potsdam and Vaihingen color map, white is the Impervious surface,red is the clutter, blue is the building, Cyan is the low vegetation, green is the tree, and yellow is the car.
+For LoveDA color map, red is the building, yellow is the road, blue is the water, purple is the barren, green is the forest, brown is the agriculture.
+Experimental setup
+Datasets and Benchmarks.
+We conduct all experiments on several widely used RS image segmentation datasets: Potsdam (Markus Gerke 2014) , Vaihingen (Markus Gerke 2014), LoveDA (Wang et al.
+2021), and iSAID (Waqas Zamir et al.
+2019).
+Following the standard MMSegmentation (Contributors 2020) configurations, we split each dataset into training and validation sets, apply cropping, and construct four semantic segmentation (SS) benchmarks: Potsdam(P), Vaihingen(V),LoveDA(L),iSAID(i), along with four domain adaptation (DA) and four domain generalization (DG) tasks: Potsdam to Vaihingen (P2V), Vaihingen to Potsdam (V2P), Rural to Urban (R2U), Urban to Rural (U2R).
+In our experimental setup, all images are cropped to a size of $512\times 512$.
+Implementation Details.
+We use MMSegmentation as our training and evaluation framework.
+We employ Mask2former (Cheng et al.
+2022) as our decoder, which is a highly efficient and effective algorithm for semantic segmentation tasks.
+For backbone, we utilize the Dinov2-Large (Oquab et al.
+2023) model to extract features, which are subsequently utilized as input for the Mask2former.
+During training, we utilize AdamW (Loshchilov and Hutter 2019) as the optimizer, with a learning rate of 1e-5 for the backbone, 1e-4 for the decoder, and 1e-4 for the relevant parameters in PEFT.
+Within the DACS training framework, we set $\lambda_{da}$ to 0.5.
+Main Result
+As depicted in Table 1, we conduct a comparative analysis of the existing mainstream PEFT methods, such as Adapter (Houlsby et al.
+2019a), LoRA (Hu et al.
+2021), and VPT (Jia et al.
+2022), against our Earth-Adapter across eight cross-domain benchmarks,with state-of-the-art method Rein (Wei et al.
+2024) in natural scene as our baseline.
+These benchmarks encompass four DA scenarios (P2V, V2P, R2U, and U2R) and four DG scenarios that mirror the DA ones.
+Figure 4: Visualization and PCA of Adapters’ Feature Maps.
+‘Agg.
+Feature’ represents the aggregated adapters’ features.
+‘PCA’ represents the Principal Component Analysis of features.
+All visualizations represent feature maps, not heatmaps.
+Thus only the semantic boundaries within the features should be focused rather than color intensities.
+Starting with the DA experiments, many existing PEFT methods suffer from severe performance degradation when compared to DA-specialized models.
+For example, LoRA achieves only 17.8% mIoU on the P2V (DA) benchmark for RS images, even lower than the Frozen DINOv2-L backbone.
+In contrast, our Earth-Adapter demonstrates much stronger domain adaptation capabilities.
+It surpasses the baseline Rein by 7.5% mIoU on P2V, 3.1% mIoU on R2U, 24.0% mIoU on U2R, and achieves an average improvement of 9.0% mIoU across all DA benchmarks.
+Shifting to the DG experiments, the issue of performance degradation remains, although the trend is somewhat mitigated.
+Unlike in the DA setting, the Frozen model highlights the potential generalizability of DINOv2-L.
+However, FFT, LoRA, and AdaptFormer continue to show clear limitations.
+Consistent with the DA benchmarks, Earth-Adapter achieves state-of-the-art (SOTA) performance in DG, surpassing the baseline Rein and Adapter by 3.1% and 4.4% mIoU, respectively.
+We also conducted comparisons on four SS benchmarks, with the results summarized in Table 2.
+Regarding PEFT-related methods, conventional approaches such as LoRA and AdaptFormer struggle to adapt to the characteristics of remote sensing images, often disrupting the representations of VFMs during fine-tuning and ultimately leading to poor performance.
+In contrast, our Earth-Adapter consistently achieves the best performance among all PEFT methods.
+We attribute this advantage to its unique design, which mitigates high-dimensional artifacts and effectively ”denoises” semantic representations, thereby enabling more accurate pixel-level feature extraction and superior semantic segmentation performance.
+Overall, our Earth-Adapter surpasses the baseline Rein by an average of 1.2% mIoU, highlighting its effectiveness.
+Ablation and Analysis
+Backbone Methods Params $\downarrow$ Train Speed (s/iter) $\downarrow$ Infer Speed (s/iter) $\downarrow$ mIoU (%) $\uparrow$
+DINOv2-L Full 304.2M 0.80 0.11 15.4
+Rein 3.0M 0.68 0.12 50.0
+Earth-Adapter 2.6M~9.6M 0.71 0.12 59.0 (+9.0)
+Table 4: Comparison on Parameters, Speed, and Performance between FFT, Rein, and Earth-Adapter.
+All results are the average score of DA benchmarks, demonstrating the excellent trade-off of Earth-Adapter.
 Adapter + HF + LF P2V (DG) V2P (DG) Avg.
@@
 The ‘+’ in the bracket represents the gain compared with one Adapter.
-Backbone Methods Params $\downarrow$ Train Speed (s/iter) $\downarrow$ Infer Speed (s/iter) $\downarrow$ mIoU (%) $\uparrow$
-DINOv2-L Full 304.2M 0.80 0.11 15.4
-Rein 3.0M 0.68 0.12 50.0
-Earth-Adapter 2.3M(1.0Msimilar-$\sim$3.9M) 0.71 0.12 59.0 (+9.0)
-Table 5: Comparison on Parameters, Speed, and Performance between FFT, Rein, and Earth-Adapter.
-All results are the average score of DA benchmarks, demonstrating the excellent trade-off of Earth-Adapter.
-Figure 2: Visualization of Predicted Segmentation Maps We Compare Earth-Adapter with the Frozen DINOv2-L backbone and our baseline Rein on eight cross-domain benchmarks.
-For the Potsdam and Vaihingen color map, white is the Impervious surface, red is the clutter, blue is the building, color is the low vegetation, green is the tree, and yellow is the car.
-For LoveDA color map, red is the building, yellow is the road, blue is the water, purple is the barren, green is the forest, brown is the agriculture.
-4.1 Experimental setup
-Datasets and Benchmark.
-We conduct all experiments on several widely used remote sensing image segmentation datasets: Potsdam [49] and Vaihingen [49], both created by the International Society for Photogrammetry and Remote Sensing (ISPRS) and LoveDA [66].
-The Potsdam dataset includes 38 very high resolution (VHR) True Orthophotos (TOP) images with 6,000 × 6,000 pixels.
-The Potsdam dataset has different imaging modes including IR-R-G channels, R-G-B channels, and R-G-B-IR channels.
-We use the RGB band for the following experiments.
-The Vaihingen dataset consists of 33 VHR TOP images with an average size of 2,494 ×2,064 pixels.
-The LoveDA dataset comprises a total of 4,190 images, each with a resolution of 1,024×1,024 pixels.
-Based on the aforementioned dataset, we follow MMSegmentation settings to split the train and val sets, crop images, and establish four benchmarks including both DA and DG: Potsdam to Vaihingen (P2V), Vaihingen to Potsdam (V2P), Rural to Urban (R2U), Urban to Rural (U2R).
-In our experimental setup, all images are cropped to a size of $512\times 512$.
-Implementation Details.
-We use MMSegmentation [11] as our training and evaluation framework.
-We employ Mask2former [10] as our decoder, which is a highly efficient and effective algorithm for semantic segmentation tasks.
-For backbone, we utilize the Dinov2 [52] series models to extract features, which are subsequently utilized as input for the Mask2former.
-During training, we utilize AdamW [47] as the optimizer, with a learning rate of 1e-5 for the backbone, 1e-4 for the decoder, and 1e-4 for the relevant parameters in PEFT.
-Within the DACS training framework, we set $\alpha$ to 0.99 and $\lambda_{uda}$ to 0.5.
-4.2 Cross-Domain Performance Comparison
-As depicted in Table 1, we conduct a comparative analysis of the existing mainstream PEFT methods, such as Adapter [28], LoRA [31], and VPT [37], against our Earth-Adapter across eight cross-domain benchmarks,with state-of-the-art method Rein [67] in natural scene as our baseline.
-These benchmarks encompass four DA scenarios (P2V, V2P, R2U, and U2R) and four DG scenarios that mirror the DA ones.
-Figure 3: Visualization and PCA of Adapters’ Feature Maps.
-‘Agg.
-Feature’ represents the aggregated adapters’ features.
-‘PCA’ represents the Principal Component Analysis of features.
-All visualizations represent feature maps, not heatmaps.
-Thus only the semantic boundaries within the features should be focused rather than color intensities.
-Figure 4: Parameter-sensitive experiments of the cut-off frequency of DFT. The results on P2V (DA) and U2R (DG) benchmarks show Earth-Adapter’s robustness to the cut-off frequency.
-Starting with the DA experiments, when compared to the performance of DA-specialized models, a significant number of existing PEFT methods exhibit performance collapse.
-For instance, LoRA achieves only 17.8% mIoU in the P2V (DA) benchmark when handling RS (Remote Sensing) images, a figure even lower than that of the Frozen backbone DINOv2-L.
-A similar trend is observed in Full Fine-Tuning, AdaptFormer, and Rein.
-We posit that this issue may stem from the DA self-training paradigm, which involves intensive image-mixing operations that these PEFT methods struggle to adapt to.
-However, two PEFT methods, Adapter and Rein, demonstrate relatively robust performance in the DA setting.
-When compared to these methods, our Earth-Adapter proves to be more effective in adapting to target domains.
-It outperforms baseline Rein by 7.5% mIoU in P2V, 3.1% mIoU in R2U, 24.0% mIoU in U2R, and 9.0% mIoU on average across the DA benchmarks.
-Notably, in the U2R (DA) benchmark, Earth-Adapter significantly mitigates the collapse of Rein, boosting the mIoU from 26.0% to 50.0%.
-Shifting to the DG experiments, the performance collapse persists, though the trend is somewhat alleviated.
-In contrast to DA, the Frozen model reveals the potential generalizability of DINOv2-L.
-Nevertheless, FFT, LoRA, and AdaptFormer still exhibit limitations.
-On the other hand VPT, and Rein continue to show strong transfer capabilities, achieving 53.5% and 53.1% mIoU on average, surpassing the Frozen model by 1.7%, and 1.3% mIoU.
-Similar to the DA benchmark, Earth-Adapter also achieves state-of-the-art (SOTA) performance in DG, with improvements of 3.1% and 4.4% mIoU over the baseline Rein and Adapter, respectively.
-4.3 Ablation Study
-Analysis on Per-class Performance.
-In this part, we demonstrate the detailed performance of each class on four benchmarks: P2V (DA), P2V (DG), U2R (DA), and U2R (DG) between Frozen DINOv2-L, Rein, and our Earth-Adapter.
-All the results are shown in Table 2, and we first discuss the P2V experiment results.
-In the P2V (DA), both Rein and Earth-Adapter show significant improvement on the Vegetation class compared with the Frozen DINOv2-L.
-Nevertheless, Earth-Adapter achieves a much higher performance than Rein, with 7.5% mIoU.
-A similar trend can be observed in the P2V (DG), where Earth-Adapter surpasses the other two methods by 7.0% and 4.1% mIoU respectively.
-Looking at the performance changes of the Vegetation and Tree classes, something interesting stands out.
-In the P2V (DA), both Earth-Adapter and Rein perform better on the Tree class than the Frozen.
-In contrast, in the P2V (DG), the performance of the Frozen with Rein on the Tree class drops drastically from 74.1% to 23.0% mIoU.
-As indicated in prior research [21], Rein’s features struggle to distinguish between the Vegetation and Tree classes due to their similar appearance from the RS perspective, especially in DG setting which faces more unpredictable domain shifts than DA.
-Moreover, in P2V (DG), comparing Rein with the Frozen, almost all other classes improve except for the Tree class.
-Thus, it can be empirically inferred that the improvement of the Vegetation class by Rein comes at the cost of a decline in the performance of the Tree class.
-However, Earth-Adapter can reduce this negative impact.
-It does this by optimizing the representation to tell the two classes apart and boost their performance.
-Figure 5: Parameter-sensitive experiments of the projector dimension in Adapter. The results on P2V (DA) and U2R (DG) benchmarks show Earth-Adapter’s robustness to dimensions.
-In the LoveDA experiments, as previously discussed, both the Frozen and Rein methods suffer from performance collapse in the U2R (DA) setting.
-When examining the performance at a more class level, it becomes evident that the primary decline for Rein occurs in the Barren and Agriculture classes.
-The performance drop in the Barren class can be attributed to it being the most sparse class within the LoveDA dataset.
-However, the performance of both the Frozen and Rein methods on the Agriculture class is a concerning 0.0% mIoU.
-In our opinion, the main factor probably is the presence of artifacts, as illustrated in Figure 1.
-Most PEFT methods from the CV field are unable to effectively handle the disturbances caused by these artifacts.
-Nevertheless, Earth-Adapter is capable of effectively utilizing the target domain data, achieving a notable 13.7% mIoU in the Barren class.
-Furthermore, as depicted in Figure 1, Earth-Adapter has a significant advantage with its frequency-domain divide-and-conquering ability, which enables it to improve the Agriculture performance from 0 to 63.2% mIoU.
-In the U2R (DG) experiments, both the Frozen and Rein methods demonstrate their effectiveness and significantly address the issues present in the DA setting.
-Additionally, Earth-Adapter achieves SOTA performance with improvements across most classes.
-Performance on Different Backbones.
-Besides the DINOv2-L backbone in Table 1, Earth-Adapter is also applied to CLIP (ViT-L) and other DINOv2 scales: ViT-Small and ViT-Base, with results in Table 3.
-These results show Earth-Adapter improves performance across diverse backbones with varying parameter scales.
-Notably, Earth-Adapter achieves larger improvements on larger models, such as ViT-L, with 7.0% and 5.7% mIoU gains, more than those on small and base scale ViTs.
-This suggests Earth-Adapter has a greater performance upper bound.
+Ablation of Different Backbones.
+Beyond the DINOv2-Large backbone used in the main results (Tables 1 and 2), we further evaluated other DINOv2 variants.
+Results for ViT-Small and ViT-Base (Table 3, rows 3–8) show that Earth-Adapter consistently improves performance across backbones of different scales.
+We also tested VFMs pre-trained on remote sensing data, including MTP-Large (Wang et al.
+2024), ScaleMAE-Large (Reed et al.
+2023), and DOFA-Large (Xiong et al.
+2024) (Table 3, rows 11–16).
+Earth-Adapter again delivers stable improvements in segmentation performance.
+Finally, we observed that DINOv2 models outperform those pre-trained on remote sensing data, likely because the latter are trained on smaller-scale datasets and thus remain less mature.
+Component Effectiveness Ablation.
+In Table 4, ablation studies on MoA’s components and adapter numbers show that using only HF or LF adapters decreases performance compared to spatial adapters, with HF causing a larger drop.
+This aligns with PCA visualizations in 2 (b) and 4, where HF features have more artifacts and noise, while LF features are smooth with clear global semantics.
+Thus, using only HF severely damages performance.
+Although only using LF features outperforms HF features (50.9%mIoU), it still underperforms spatial features (56.9% mIoU) due to lacking semantic details.
+After that, we also conduct more experiments on the number of Spatial adapters, and reveal the best composition is the one spatial adapter with one HF and LF adapters.
 Prediction Visualization Analysis.
@@
 And in other experiments, Earth-Adapter also exhibits better prediction than Rein and Frozen backbone, showing a higher performance upperbound.
+Method P2V(DG) V2P(DG) Avg.
+w/o DR 63.6 53.2 58.4
+w/ DR 64.9 55.1 60.0(+1.6)
+Table 6: Ablation of Dynamic Router.
+Feature Visualization Analysis.
+In Figure 4, we visualize features captured by three adapters on P2V (DG) and R2U (DG) benchmarks.
+As discussed in the introduction, LF features focus on coarse-grained and global semantics, while HF features show detailed representations.
+The aggregated features are a weighted summation of the three different frequency features.
+PCA visualizations clearly show each feature’s characteristics.
+Take P2V in the figure as an example, artifacts are almost filtered into high-frequency features (shown in ‘HF PCA’), and dynamic fusion of all features ensures the final aggregated features’ PCA maintains clear semantic edges and successfully filters out artifacts.
+These visualizations further enhance Earth-Adapters’ explainability.
+Performance-Speed Trade-Off Analysis.
+We compare the parameters, speed, and performance between FFT, Rein, and Earth-Adapter in Table 4.
+Significantly, at the same time of close training and inference speed with Rein, Earth-Adapter achieves better performance (9.0% mIoU improvement), with a smaller parameter scale.
+This further confirms the efficiency of Earth-Adapter, which can be attributed to its simple yet effective design methodology, making it particularly well-suited for semantic segmentation tasks on remote sensing images.
+Ablation of Dynamic Router.
+As shown in Tab.
+6, the use of the dynamic router leads to better results, yielding an average improvement of 1.6% mIoU across the two DG benchmarks.
+Compared with static weights (where we assign a weight of 1/3 to each static route), the dynamic router can adaptively adjust the feature allocation weights, enabling more effective representation optimization.
+Conclusion
+Vision Foundation Models (VFMs) excel across diverse visual tasks, and pairing them with Parameter-Efficient Fine-Tuning (PEFT) is an effective way to adapt them to downstream applications.
+However, existing PEFT methods often fall short on remote sensing (RS) semantic segmentation because they fail to suppress the pervasive artifacts in VFM deep features for RS imagery.
+To address this limitation, we introduce Earth-Adapter, a simple yet effective PEFT framework tailored for RS segmentation.
+It employs a frequency-guided mixture of adapters (MoA) that isolates artifacts into the high-frequency subspace, fine-tunes features within each subspace, and adaptively fuses them through a learnable router.
+Acknowledgments
+This work was partly supported by National Natural Science Foundation of China (62301046, 62506229) and Natural Science Foundation of Shanghai (25ZR1402268).
+Appendix
+Training Details
+Dataset and Benchmark
+Dataset Resolution Categories Train split Val Split
+Potsdam $6000\times 6000$ 6 24 14
+Vaihingen $\sim 2494\times 2064$ 6 16 17
+LoveDA $1024\times 1024$ 7 2521 1668
+iSAID $800\times 800\sim 13000\times 13000$ 16 1403 468
+Table 7: Dataset overview
+Spatial Adapter MoA Frequency Layer Cutoff Frequency U2R (DA) V2P (DA) P2V (DG) R2U (DG)
+- - - 48.6 60.1 61.0 58.6
+Full 0.3 47.0 60.9 61.3 58.1
+shallow 3 0.3 48.3 60.1 63.8 58.0
+shallow 6 0.3 47.6 60.1 63.8 59.0
+deep 3 0.3 47.4 61.6 63.7 58.7
+deep 6 0.3 48.0 60.6 64.9 58.5
+deep 3 / deep 6 0.2 / 0.3 50.0 62.2 64.9 59.0
+Table 8: Ablation of Frequency Adapter with different activation layers.
+The results demonstrate that the Earth-Adapter tends to be more beneficial when used in deeper network layers.
+The results in the gray shadow mean the best performance using deep 3 or 6 layers and 0.2 or 0.3 cutoff frequencies.
+Spatial Adapter MoA Frequency Layer Cutoff Frequency U2R (DA) V2P (DA) P2V (DG) R2U (DG)
+- - - 48.6 60.1 61.0 58.6
+Full 0.3 47.0 60.9 61.3 58.1
+shallow 3 0.3 48.3 60.1 63.8 58.0
+shallow 6 0.3 47.6 60.1 63.8 59.0
+deep 3 0.3 47.4 61.6 63.7 58.7
+deep 6 0.3 48.0 60.6 64.9 58.5
+deep 3 / deep 6 0.2 / 0.3 50.0 62.2 64.9 59.0
+Table 9: Ablation of Frequency Adapter with different activation layers.
+The results demonstrate that the Earth-Adapter tends to be more beneficial when used in deeper network layers.
+The results in the gray shadow mean the best performance using deep 3 or 6 layers and 0.2 or 0.3 cutoff frequencies.
+Lr 1e-4
+Training iters 20k/40k
+Batch size 4/8
+$\alpha$ 0.99
+$\lambda_{da}$ 0.5
+Adapter dim 16/24/32/64
+Cutoff frequency 0.2/0.3
+Activation layer pre3/pre6/suf3/suf6
+Table 10: Hyper-parameter overview
+Layer Adapter Earth-Adapter
+U2R (DA) R2U (DG) U2R (DA) R2U (DG)
+Frozen 11.8 57.1 11.8 57.1
+[0,1,2,3,4,5] 44.7 57.4 47.6 59.0
+[18,19,20,21,22,23] 46.0 58.1 50.0 58.5
+Table 11: Layer effect comparison between Adapter and Earth-Adapter. The results show Earth-Adapter learns better under the same layer setting.
 Benchmark Adapter dim Cutoff frequency Activation layer mIoU(%)
@@
 All the results are the best performance.
-Feature Visualization Analysis.
-In Figure 3, we visualize features captured by three adapters on P2V (DG) and R2U (DG) benchmarks.
-As discussed in the introduction, LF features focus on coarse-grained and global semantics, while HF features show detailed representations.
-The aggregated features are a weighted summation of the three different frequency features.
-PCA visualizations clearly show each feature’s characteristics.
-Take P2V in the figure as an example, artifacts are almost filtered into high-frequency features (shown in ‘HF PCA’), and dynamic fusion of all features ensures the final aggregated features’ PCA maintains clear semantic edges and successfully filters out artifacts.
-These visualizations further enhance Earth-Adapters’ explainability.
-Component Effectiveness Ablation.
-In Table 4, ablation studies on MoA’s components and adapter numbers show that using only HF or LF adapters decreases performance compared to spatial adapters, with HF causing a larger drop.
-This aligns with PCA visualizations in 1 (b) and 3, where HF features have more artifacts and noise, while LF features are smooth with clear global semantics.
-Thus, using only HF severely damages performance.
-Although only using LF features outperforms HF features (50.9%mIoU), it still underperforms spatial features (56.9% mIoU) due to lacking semantic details.
-After that, we also conduct more experiments on the number of Spatial adapters, and reveal the best composition is the one spatial adapter with one HF and LH adapters.
-Performance-Speed Trade-Off Ablation.
-We compare the parameters, speed, and performance between FFT, Rein, and Earth-Adapter in Table 5.
-Significantly, at the same time of close training and inference speed with Rein, Earth-Adapter achieves better performance (9.0% mIoU improvement), with a smaller parameter scale.
-Hyper-Parameter Sensitive Ablation.
-In this section, we conduct an ablation study to test Earth-Adapter with two hyper-parameter variations on both the P2V (DA) and U2R (DG) benchmarks.
-Figure 4 presents the studies of the cut-off frequency with the dimension as 16, and Figure 5 illustrates the studies of the projector dimension in the Adapter with the cut-off frequency as 0.3.
-Both experiments show a consistent trend: Earth-Adapter shows greater robustness in the DA than in the DG setting.
-We posit that this occurs because the DA task inherently possesses higher certainty compared to DG.
-In the DA, even when hyper-parameters are altered, Earth-Adapter can effectively learn from the augmented target domain data, thereby maintaining performance stability.
-Despite DG’s uncertainty, Earth-Adapter shows a relatively stable performance in a certain parameter range.
-For instance, when the cut-off frequency is between 0.3 and 0.4, performance fluctuations are minimal.
-Similarly, for the projector dimension, performance is stable across a broad range from 16 to 256 and only fluctuates when the dimension is 512.
-Importantly, there are no other hyper-parameters requiring tuning.
-These results underscore Earth-Adapter’s superiority in handling parameter variations, offering a reliable solution for both DA and DG.
-Layer Adapter Earth-Adapter
-U2R (DA) R2U (DG) U2R (DA) R2U (DG)
-Frozen 11.8 57.1 11.8 57.1
-[0,1,2,3,4,5] 44.7 57.4 47.6 59.0
-[18,19,20,21,22,23] 46.0 58.1 50.0 58.5
-Table 7: Layer effect comparison between Adapter and Earth-Adapter. The results show Earth-Adapter learns better under the same layer setting.
-Spatial Adapter MoA Frequency Layer Cutoff Frequency U2R (DA) V2P (DA) P2V (DG) R2U (DG)
-- - - 48.6 60.1 61.0 58.6
-Full 0.3 47.0 60.9 61.3 58.1
-shallow 3 0.3 48.3 60.1 63.8 58.0
-shallow 6 0.3 47.6 60.1 63.8 59.0
-deep 3 0.3 47.4 61.6 63.7 58.7
-deep 6 0.3 48.0 60.6 64.9 58.5
-deep 3 / deep 6 0.2 / 0.3 50.0 62.2 64.9 59.0
-Table 8: Ablation of Frequency Adapter with different activation layers.
-The results demonstrate that the Earth-Adapter tends to be more beneficial when used in deeper network layers.
-The results in the gray shadow mean the best performance using deep 3 or 6 layers and 0.2 or 0.3 cutoff frequencies.
-4.4 Analysis of Overall Parameter Configuration.
+We conduct experiments on several major optical remote sensing image segmentation datasets.
+The original datasets have varying resolutions, ranging from $1024\times 1024$ to $6000\times 6000$, which we uniformly crop to $512\times 512$ for training and evaluation.
+The detailed information of the datasets are shown in the table below.
+Training Framework
+In this study, we explore the application of Earth-Adapter across three distinct tasks: In-Domain(SS), DG and DA Semantic Segmentation.
+For SS and DG tasks, we train Earth-Adapter with Mask2former in end-to-end supervised learning as Equation 2.
+For DA Semantic Segmentation, we use the self-training framework introduced in DACS (Tranheden et al.
+2021) without using target domain labels.
+The pseudo labels for target domain images are produced by a teacher network $G_{\theta,\xi,\epsilon}^{\dagger}$:
+$\mathbf{\bar{y}_{T}}^{(j,k)}=[k=\mathop{\arg\max}_{k^{\prime}}G_{\theta,\xi,\epsilon}^{\dagger}(\mathbf{x_{T}})))^{(j,k^{\prime})}],$ (14)
+where $\cdot$] denotes the Iverson bracket, and $j\in HW$ is the spatial index of the pixel in the image.
+The teacher network remains inactive during the training process, with its parameters being updated based on an exponential moving average (EMA) of the student model’s parameters:
+$G_{\theta,\xi,\epsilon,t+1}^{\dagger}=\alpha G_{\theta,\xi,\epsilon,t}^{\dagger}+(1-\alpha)G_{\theta,\xi,\epsilon,t}$ (15)
+Here $t$ denotes the $t_{th}$ training iteration.
+The hyperparameter $\alpha$ controls the update speed of the teacher network’s parameters.
+After obtaining the pseudo-labels, we perform category-wise mixing of samples from the source and target domains:
+$\mathbf{x}_{mix}$ $=\mathbf{x_{S}}\odot\mathcal{M}+\mathbf{x_{T}}\odot(1-\mathcal{M})$ (16)
+$\mathbf{y}_{mix}$ $=\mathbf{y_{S}}\odot\mathcal{M}+\mathbf{\bar{y}_{T}}\odot(1-\mathcal{M})$
+where $(\mathbf{x_{S}},\mathbf{y_{S}})$ is the data sample from the source domain, and $\mathcal{M}$ is generated by randomly selecting half of the categories based on the labels of the source domain $\mathbf{y_{S}}$.
+The Optimization objective for the DA task if defined by:
+$\mathop{\arg\min}_{\theta,\epsilon,\xi}$ $\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}\cup\mathcal{D}_{mix}}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}),\mathbf{y}\right)$ (17)
+$\quad+\lambda_{uda}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}_{mix}),\mathbf{y}_{mix}\right).$
+Hyper-Parameter Configuration
+As shown in Table 7, we provide an overview of training hyperparameters.
+Specifically, for DA and DG tasks, we train for 20k iterations with a batch size of 4; for SS tasks, we train for 40k iterations with a batch size of 8.
+Notably, for the iSAID SS task, we train for 80k iterations with a batch size of 8.
+Other configurations are detailed in rows 5 to 7 of the table.
+More Ablation and Analysis
+Analysis of Overall Parameter Configuration.
 The detailed hyperparameter configurations for each benchmark are presented in Table 12.
-Overall, our method demonstrates exceptional robustness, exhibiting relatively low sensitivity to parameter variations.
+Overall, our algorithm demonstrates exceptional robustness, exhibiting relatively low sensitivity to parameter variations.
 Below we provide an in-depth analysis of the hyperparameter effects.
@@
 Our study primarily focused on the Dinov2-Large model, a 24-layer Vision Transformer architecture.
-We evaluated four distinct adapter configurations: {0,1,2}, {0,1,2,3,4,5}, {21,22,23}, and {18,19,20,21,22,23}.
+We evaluated four distinct adapter configurations: [0,1,2], [0,1,2,3,4,5], [21,22,23], and [18,19,20,21,22,23].
 As shown in Table 12, different benchmarks achieved optimal performance with varying frequency-domain adapter configuration.
-For instance, P2V (DA) attained its peak performance of 67.7% mIoU with adapters in 0,1,2, while P2V (DG) achieved its best results with the {18,19,20,21,22,23} configuration.
+For instance, P2V (DA) attained its peak performance of 67.7% mIoU with adapters in [0,1,2], while P2V (DG) achieved its best results with the [18,19,20,21,22,23] configuration.
 Similar to the adapter dimension findings, the frequency adapter layers demonstrate task-dependent optimization patterns.
-Notably, three DG tasks (P2V (DG), V2P (DG), and U2R (DG)) consistently performed best with adapters in the layers {18,19,20,21,22,23}, suggesting that domain generalization (DG) requires more adapter parameters to capture general semantic representations to improve cross-domain performance.
-4.5 Discussion on Adapter Layer
+Notably, three DG tasks (P2V (DG), V2P (DG), and U2R (DG)) consistently performed best with adapters in the layers [18,19,20,21,22,23], suggesting that domain generalization (DG) requires more adapter parameters to capture general semantic representations to improve cross-domain performance.
 Layer-wise comparison between Adapter and Earth-Adapter.
 We analyze the layer-wise behavior of Adapter and compare the performance of Adapter and Earth-Adapter under different layer configurations (Table 11).
-While prior work [71] suggests that deeper layers capture more semantically meaningful features, our Earth-Adapter demonstrates superior performance in both shallow-layer (first six) and deep-layer (last six) fine-tuning.
+While prior work (Oquab et al.
+2023) suggests that deeper layers capture more semantically meaningful features, our Earth-Adapter demonstrates superior performance in both shallow-layer (first six) and deep-layer (last six) fine-tuning.
 For instance, in the U2R (DA) task, Earth-Adapter (shallow fine-tuning) outperforms the frozen model by 35.8% mIoU, while in R2U (DG), it surpasses the frozen baseline by 1.9% mIoU.
@@
 Our experiments reveal that these frequency layers play a crucial role in the overall performance of Earth-Adapter.
-Notably, integrating the Frequency Adapter results in a slight performance drop in two benchmarks: U2R (DA) ($-1.6\%$ mIoU) and R2U (DG) ($-0.5\%$ mIoU).
-However, with an efficient parameter search, Earth-Adapter still surpasses the Spatial Adapter.
+Notably, integrating the Frequency Adapter results in a slight performance drop in two benchmarks: U2R (DA) ($1.6\%$ mIoU) and R2U (DG) ($0.5\%$ mIoU).
+However, with an efficient parameter search, Earth-Adapter still surpasses the Spatial-Adapter.
 As shown in the table, Earth-Adapter achieves 50.0% mIoU in the U2R (DA) task, outperforming Spatial-Adapter by 1.4% mIoU.
-Similarly, it attains 59.0% mIoU in the R2U (DG) task, exceeding Spatial Adapter by 0.4% mIoU.
-A key observation is that Earth-Adapter exhibits sensitivity to frequency layer configurations and cutoff frequencies under certain benchmarks—an aspect we aim to refine in future work.
-5 Conclusion
-In this paper, we first highlight the performance drop when transferring existing PEFT methods to the RS cross-domain experiments.
-Detailedly, we analyze these problems through the research [12] and our artifacts-related findings, then raise that existing methods can not effectively handle the artifacts appearing in RS images.
-To tackle this, we introduce Earth-Adapter, our RS-specialized PEFT method.
-It is a simple-yet-effective and the first PEFT method for cross-domain RS semantic segmentation.
-In the experiment section, we first comprehensively evaluate Earth-Adapter’s performance and compare with other methods on widely-used benchmarks.
-Secondly, we analyze Earth-Adapter’s effectiveness with different backbones, as well as the trade-off between parameters, speed, and performance.
-Finally, we visualize the decoder’s predictions and expert outputs.
-Our results demonstrate Earth-Adapter’s power, efficiency, robustness, and explainability.
-However, there is still space for improvement.
-For example, it’s still unclear on which backbone layers it should be applied, and how to further enhance the utilization of frequency features.
-In our subsequent work, we’ll pursue more potential optimization.
-The code will be public soon, and we hope Earth-Adapter will be helpful to this community.
+Similarly, it attains 59.0% mIoU in the R2U (DG) task, exceeding Spatial-Adapter by 0.4% mIoU.
+A key observation is that Earth-Adapter exhibits sensitivity to frequency layer configurations and cutoff frequencies under certain benchmarks—which we aim to refine in future work.